### v1

# Parameter-Efficient Augment Plugin for Class-Incremental Learning

Abstract
Existing class-incremental learning (CIL) approaches based on replay or knowledge distillation are often constrained by forgetting or the stability-plasticity dilemma. Some expansion-based approaches could achieve higher accuracy. However, they always require significant parameter increases. In this paper, we propose a plugin extension paradigm termed the Deployment of extra LoRA Components (DLC) for non-pre-trained CIL scenarios. We treat the feature extractor trained through replay or distillation as a base model with rich knowledge. For each task, we use Low-Rank Adaptation (LoRA) to inject task-specific residuals into the base model’s deep layers. During inference, representations with task-specific residuals are aggregated to produce classification predictions. To mitigate interference from non-target LoRA plugins, we introduce a lightweight weighting unit. This unit learns to assign importance scores to different LoRA-tuned representations. Like downloadable contents in software, our method serves as a plug-and-play enhancement that efficiently extends the base methods. Remarkably, on the large-scale ImageNet‑100, with merely 4% of the parameters of a standard ResNet-18, our DLC model achieves a significant 8% improvement in accuracy, demonstrating exceptional efficiency. Moreover, it could surpass state-of-the-art methods under the fixed memory budget.
1 Introduction
In recent years, deep learning has achieved remarkable success across a variety of domains [jaimes2007multimodal, ye2019learning, chen2021large, chen2022learning, yang2024entaugment]. However, most models rely on a key assumption that the data distribution and learned classes remain fixed. In real-world scenarios, data distributions evolve continuously and new classes emerge over time [gomes2017survey], making static training paradigms impractical. Moreover, learning each new task typically relies on introducing additional task-specific sub-models, leading to inefficient model growth over time. Class-Incremental Learning (CIL) [zhou2024class, liangloss, yang2025supervised] thus plays a crucial role in enabling models to continually learn from non-stationary data streams while retaining prior knowledge.
Existing CIL methods  [zhou2024class] can be broadly categorized into rehearsal-based, restriction-based, and expansion-based approaches. Rehearsal-based methods maintain a small replay buffer by storing a limited number of representative exemplars from previous tasks [lopez2017gradient, rolnick2019experience, de2021rep]. However, the number of stored exemplars is typically far smaller than that of samples in the current task, creating a severe data imbalance that biases learning toward new classes and thus exacerbates catastrophic forgetting. Restriction-based methods [nguyen2017variational, ahn2019uncertainty] aim to mitigate catastrophic forgetting by constraining model updates during new task learning. A representative example is knowledge distillation [rebuffi2017icarl, simon2021learning], which preserves previous knowledge by aligning the current model’s output with those of its earlier versions. Such constraints introduce a trade-off: the model must balance adherence to past knowledge with adaptation to new data. This stability-plasticity dilemma fundamentally limits the achievable performance of restriction-based approaches.
Figure 1: Parameter-accuracy comparison of different CIL methods on ImageNet-100 with ResNet-18. DLC achieves the best balance between parameter efficiency and accuracy compared to prior SOTA methods. iCaRL-DLC, WA-DLC, and BiC-DLC mean integrating DLC into the corresponding methods, respectively.
To address this, expansion-based CIL methods [mallya2018piggyback, serra2018overcoming, xu2018reinforced] allocate task-specific architectural blocks, creating isolated parameter spaces for new knowledge and freezing previously learned parameters. While this isolates interference between tasks, the scalability limits as each new task requires its own feature extractor, causing parameters to grow linearly and incurring heavy storage costs. Recent pre-trained models-based methods [zhou2024continual] attempt to reduce this cost through lightweight fine-tuning modules. Yet, most PTM-based designs rely heavily on vision transformers with massive parameters (86M), and their fine-tuning modules can even surpass the size of conventional models. For instance, the introduction of a single Adapter module [zhou2024revisiting] incurs an parameter overhead (15M) greater than that of a complete ResNet-18 (11M). In addition, these PTM-based expansion methods face practical limitations, as PTMs are not universally available in practical applications. Therefore, expansion-based methods raise concerns despite strong performance. Their long-term sustainability and scalability in lifelong learning are questionable. This exposes a new quandary to CIL: How can we maintain high accuracy without a prohibitive linear increase in parameters as the number of tasks grows?
To address this challenge, we propose the Deployment of extra LoRA Components (DLC). Rather than expanding the feature extractor model, DLC adopts a plug-and-play extension paradigm. Specifically, we attach task-specific Low-Rank Adaptation (LoRA) plugins to the deep layers of the base model. These plugins are updated exclusively during their corresponding task’s training phase, which occurs after the base model has been trained. This design ensures that the DLC plugins are trained independently, thus enabling their seamless and non-interfering integration into other methods. During inference, the base model loads all LoRA plugins to extract their feature representations, and they are aggregated into a composite one for final classification. However, non-target LoRA plugins that were not adequately updated on the dataset corresponding to the sample’s task may introduce noise or less informative residuals into the composite representation. So we propose a dedicated lightweight weighting unit before the classifier. This unit automatically assigns importance weights to different LoRA plugin outputs. DLC could enhance many non-expansion CIL approaches, such as knowledge distillation or replay methods. Experimental results on various benchmark datasets show that our method consistently improves the performance of replay- and distillation-based baselines. For instance, it could achieve up to 8% accuracy gains on ImageNet-100 using ResNet-18 and on CIFAR-100 using ResNet-32, while only increasing the total number of parameters by 4% and 11%, respectively. Furthermore, under the same memory budget, DLC outperforms existing state-of-the-art baselines. Our contributions are summarized as follows:
  • We propose a plugin extension framework DLC, which can be seamlessly integrated into diverse CIL methods without architectural modification by injecting task-specific residuals into deep layers via LoRA.
  • We introduce a lightweight weighting unit for inference, which adaptively suppresses irrelevant LoRA activations according to their learned importance. This adaptive filtering mechanism effectively mitigates feature interference between old and new tasks, leading to more stable and accurate incremental classification.
  • Extensive experiments show that DLC significantly boosts various CIL methods’ performance. Meanwhile, under the fixed memory budget, the DLC-enhanced distillation method could surpass SOTA baselines.
2 Related Work
DLC can be viewed as a specialized form of expansion-based CIL that augments base methods with task-specific plugins. This section first introduces methods facing the stability-plasticity dilemma that are compatible with DLC, followed by a review of expansion approaches.
2.1 CIL with stability-plasticity dilemma
To mitigate catastrophic forgetting, many methods implement constrained parameter updates during model optimization [nguyen2017variational, bian2024make, cui2025few]. A prominent strategy in this category is knowledge distillation [simon2021learning, huang2024etag]. It protects the previously acquired knowledge through a modified distillation loss from the old to the new model. iCaRL [rebuffi2017icarl] mitigates forgetting by distilling the logits from the old model into the new model. Subsequent methods like BiC [wu2019large] and WA [zhao2020maintaining] further propose bias correction and weight alignment techniques to calibrate the classifier. Moreover, these methods preserve extra knowledge of prior classes by employing replay buffers [shaul2024teal, aghasanli2025prototype] to save exemplars in training. Furthermore, recent research has developed plug-and-play auxiliary loss functions [gao2025maintaining] that can be readily integrated into existing methods to enhance their accuracy. However, the fixed capacity of the feature extractor poses a fundamental constraint. The model is inherently limited in its ability to retain knowledge from past tasks and learn new tasks effectively, forcing it to seek a balance between stability and plasticity through a trade-off between classification and distillation/regularization losses, which ultimately caps its achievable accuracy.
2.2 Expansion-Based CIL
Such methods expand specialized parameters for model [mallya2018piggyback, serra2018overcoming, xu2018reinforced]. DER and TagFex [yan2021dynamically, zheng2025task] train an independent representation model on the new task for feature expansion, while the models of old tasks are frozen to retain the learned knowledge. MEMO [zhou2022model] splits the feature extractor into a generic shallow module and several task-specific deep modules. By increasing task-specific modules with fewer parameters, MEMO could save more memory space for storing replay samples than model-level expansion methods. Expansion-based methods achieve notably higher accuracy ceilings compared to fixed-capacity models. However, these methods result in a linear increase in the total number of parameters with tasks, imposing a significant storage overhead. Specifically, in the case of TagFex, the introduced attention module further increases the parameter overhead of the expansion. Even MEMO still expands over 70% of the extractor’s parameters for every new task. Emerging research leverages pre-trained models (PTMs) [zhou2024revisiting, mcdonnell2024ranpac, tan2024semantically] for CIL, mitigating catastrophic forgetting by leveraging knowledge encapsulated in PTMs. This has spurred the development of specialized parameter expansion methods designed for PTMs, which involve training lightweight fine-tuning modules like adapters [zhou2024expandable, sun2025mos]. However, such methods paradoxically result in higher overall parameter counts rather than a reduction. This is primarily due to the large parameter size of ViT backbones—significantly exceeding that of ResNet—combined with the non-negligible overhead of their tuning modules. An adapter introduces parameters equivalent to 15% of the base ViT’s total. Furthermore, PTM is not available in many CIL cases. These parameter-efficient module expansion methods depend on the rich prior knowledge embedded in PTMs and cannot be directly transferred to non-pre-trained models.
Figure 2: The proposed DLC framework. Left: Training. When a new task $t$ arrives, a dedicated plugin set $L_{t}$ is created. Then sequentially train the feature extractor and $L_{t}$. Right: Test. The feature extractor sequentially loads all task plugins $L_{t}$ to produce enhanced representations, which are concatenated, weighted by a gating unit, and classified.
3 Preliminaries
3.1 Problem Formulation of CIL
A CIL model aims to learn new classes while retaining previous knowledge from the continuously evolving data streams $D_{1},D_{2},\cdots,D_{t}$ [de2021continual], where $D_{i}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{n_{i}}$ containing $n_{i}$ instances. One class will appear only once on a given $D_{i}$ and will not be repeated. At each incremental learning stage $t$, only the current data stream $D_{t}$ and a few exemplars from the previous classes [rebuffi2017icarl] can be used for model training. For any stage $t$, the model has learned $D=D_{1}\cup D_{2}\cup\cdots\cup D_{t}$ streams, We aim to find a model $f(\mathbf{x}):\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes empirical risk across all learned classes, where $\mathcal{X}$ and $\mathcal{Y}=Y_{1}\cup Y_{2}\cdots\cup Y_{t}$ represent the sample’s feature and class spaces, respectively. The effectiveness of the model is evaluated by the test dataset $D^{test}=D_{1}^{test}\cup D_{2}^{test}\cup\cdots\cup D_{t}^{test}$, which can be expressed in Eq.(1):
$f^{*}(\mathbf{x})=\arg\min_{f\in\mathbb{H}}\mathop{\mathbb{E}}\nolimits_{(\mathbf{x},y)\in D^{test}}[\mathbb{I}(f(\mathbf{x})\neq y)],$ (1)
where $\mathbb{H}$ is the hypothesis space of model and $\mathbb{I}(\cdot)$ denotes the indicator function.
3.2 –
Consistent Knowledge Preserved in CIL Considering the classic CIL scenario with only a single feature extractor. The objective function used in non-CIL settings is the cross-entropy loss $\mathcal{L}_{\text{CE}}$, while restriction-based CIL methods incorporate an additional restriction term $\mathcal{L}_{\text{MEM}}$ that enforces the preservation of knowledge acquired from previous tasks. This composite objective function can be formally expressed as:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{MEM}}.$ (2)
As a canonical example of restriction-based methods, we examine logit-based distillation, which ensures consistency between the logit distributions produced by the old and new models for known classes. Specifically, the restriction term $\mathcal{L}_{\text{MEM}}$ is defined as a distillation loss $\mathcal{L}_{\text{KD}}$. Let $\hat{q}(\boldsymbol{x})$ and $q(\boldsymbol{x})$ represent the logits of the teacher and student models. The cross-entropy version of $\mathcal{L}_{\text{KD}}$ is defined as:
$\mathcal{L}_{\text{KD}}=-\sum_{i}\hat{q}_{i}(\boldsymbol{x})\log(q_{i}(\boldsymbol{x})),$ (3)
and the KL divergence version is defined as:
$\mathcal{L}_{\text{KD}}=\tau^{2}\sum_{i}\hat{q}_{i}(\boldsymbol{x})\log(\frac{\hat{q}_{i}(\boldsymbol{x})}{q_{i}(\boldsymbol{x})}),$ (4)
where the $\tau$ is the temperature scalar.
Moreover, most methods are complemented by a replay buffer that stores a subset of representative exemplars from previous tasks. These exemplars are interleaved with data from the current task during training to help maintain performance on past tasks. Specifically, when learning the $t$-th task, the model is conventionally trained only on the new dataset $D_{t}$. With the replay buffer, the training dataset is supplemented with $D_{\text{exp}}$ retrieved from the buffer, forming a combined training set $D_{\text{train}}=D_{t}\cup D_{\text{exp}}$.
Although the model’s parameters are dynamically updated throughout the incremental learning process, the use of knowledge distillation and the replay buffer ensures that a portion of knowledge from previous tasks remains preserved. Replay ensures the model continues to minimize the empirical risk on a representative subset of past tasks, anchoring the parameters in regions consistent with previous data. Concurrently, knowledge distillation functions as a functional regularizer, ensuring that the mappings for old classes remain stable by aligning the output logits (or distributions) with those of the prior model. Consequently, a model trained with the replay buffer and distillation retains significant knowledge about past classes, functioning as a capable but suboptimal ”generalist.” We posit that the knowledge stored in the model can be efficiently activated and tuned with minor task-specific parameters.
4 The Proposed Method
While the use of rehearsal and restriction enable the model to retain knowledge from previous tasks, an additional objective or imbalanced exemplars could create an optimization trade-off. This results in a suboptimal convergence point for the cross-entropy loss or past data, ultimately capping the model’s classification performance.
DLC proposes a distinctive paradigm, as shown in Fig. 2. We regard the model trained with a replay buffer and distillation as a foundation model, which stored knowledge across all tasks (albeit with suboptimal accuracy). Then refine it through residuals from each task-specific LoRA plugin to enhance accuracy. The details of the proposed DLC framework are elaborated in the following subsections.
4.1 LoRA for Convolutional Layers
A basic CIL model typically comprises a feature extractor $\phi(\boldsymbol{x})$ followed by a classifier $\boldsymbol{W}$. Most model expansion methods focus on expanding the $\phi(\boldsymbol{x})$ to achieve feature expansion. To maximize parameter efficiency, our approach avoids expanding the entire $\phi(\boldsymbol{x})$ or its deeper layers. Instead, we append parameter-efficient plugins after selected network layers within the $\phi(\boldsymbol{x})$ to adapt the representations produced by the base model. We specifically adopt Low-Rank Adaptation (LoRA) [hu2022lora] due to its minimal parameter size and its architectural flexibility.
The standard LoRA method was originally designed for linear layers. Given that most non-pre-trained CIL models are based on Convolutional Neural Networks (CNNs), it is necessary to change the original LoRA formulation to the convolutional operations fundamental to CNNs. For a convolutional layer with weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times K\times K}$ and input $\boldsymbol{x}\in\mathbb{R}^{N\times C_{\text{in}}\times H\times W}$, the original forward pass is $\boldsymbol{h}=W_{\text{conv}}\ast\boldsymbol{x}$, where $\ast$ denotes the convolution operation. By analogy to LoRA, ConvLoRA constructs a low-rank adaptation for a convolutional weight tensor via a sequence of two convolutional operations. The forward pass first processes the input with $\mathbf{A}$:
$\boldsymbol{z}=\mathbf{A}\ast\boldsymbol{x},$ (5)
where $\mathbf{A}\in\mathbb{R}^{r\times C_{\text{in}}\times K\times K}$ projects the feature to a low-rank space. Then followed by kernel B:
$\boldsymbol{h}_{\text{lora}}=\mathbf{B}\ast\boldsymbol{z},$ (6)
where $B\in\mathbb{R}^{C_{\text{out}}\times r\times 1\times 1}$ restores channel dimensionality. In summary, the forward pass with ConvLoRA is:
$\boldsymbol{h}^{\prime}=W_{\text{conv}}\ast\boldsymbol{x}+\frac{\alpha}{r}(B\ast(A\ast\boldsymbol{x})),$ (7)
where $\alpha$ is a scaling hyperparameter, $r$ is the low-rank.
4.2 Deploy LoRA as Task-Specific Plugins
LoRA cannot be directly applied in non-PTM settings, not only because PTMs provide diverse knowledge. Without a frozen PTM backbone, the output of the model changes with parameter updates, leading to unstable layer inputs for LoRA. However, LoRA becomes applicable in replay- and distillation-based methods, as these approaches effectively constrain the variation of the model’s outputs on old tasks, thereby ensuring that the input features to LoRA remain close to the distribution encountered during its training phase. Specifically, after training on task $t$, the output deviation of the model at any layer $\ell$ will not exceed $K_{\ell}\Gamma_{t}$ (The detailed proof is provided in the supplementary material). Here, $K_{\ell}$ is determined by hyperparameters such as the network structure and distillation temperature, while $\Gamma_{t}$ characterizes the intensity of distribution discrepancy induced by selected distillation and replay strategies. This upper bound limits the maximum feature drift magnitude of the model on the old tasks’ distribution.
Similar to expansion-based methods, our approach incorporates independent LoRA plugins for each task. Specifically, for each task $t$, we create a dedicated set of $k$ LoRA plugins for $\phi(\boldsymbol{x})$, denoted collectively as $L_{t}$. The $\phi(\boldsymbol{x})$ equipped with $L_{t}$ is denoted by $\phi^{L_{t}}(\boldsymbol{x})$. In DLC, LoRA plugins are deployed in the final layers of $\phi(\boldsymbol{x})$. Taking a CNN-based architecture as an example, the first LoRA plugin in $L_{t}$ is applied to the last convolutional layer of the $\phi(\boldsymbol{x})$. The remaining LoRA plugins are then sequentially deployed in a backward manner toward earlier layers. This design is motivated by two key considerations. First, features extracted from deeper layers are more semantically meaningful and are directly utilized for final classification decisions. Injecting task-specific residual adaptations via LoRA at this level allows for more direct enhancement of inter-class discriminability. Second, it demonstrates better compatibility with distillation-based objectives. When applied to shallow layers, the residual modifications introduced by LoRA would cause alterations in low-level features that propagate through the entire network, consequently amplifying the divergence in the final output logits. This increased discrepancy substantially raises the optimization difficulty for distillation losses, which rely on logit alignment between old and new models.
In DLC, the LoRA plugins and $\phi(\boldsymbol{x})$ undergo a decoupled, two-phase training procedure. Specifically, upon the arrival of a new task $t$, we first train only the $\phi(\boldsymbol{x})$ and classifier $\boldsymbol{W}$, keeping all LoRA plugins frozen, to facilitate adaptation to the new task distribution. After this phase, $\phi(\boldsymbol{x})$ is frozen to prevent knowledge drift, and the $L_{t}$ are trained with the same data. Upon completion, $L_{t}$ is permanently frozen to safeguard the acquired task-specific knowledge from forgetting. The training objective for all LoRA plugins incorporates the cross-entropy loss and the auxiliary loss [yan2021dynamically, zhou2022model] commonly employed in expansion-based methods, formulated as follows:
$\mathcal{L}_{\text{LoRA}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}.$ (8)
The two-phase training procedure ensures that the LoRA plugins are trained without interference from distillation losses, enabling specialized learning for the current task and thereby improving its accuracy. Since the $\phi(\boldsymbol{x})$ remain frozen during LoRA updates, the training process of LoRA plugins exhibits stability. Furthermore, it prevents the $\phi(\boldsymbol{x})$’s training from being affected by LoRA, thereby maintaining the plug-and-play characteristic of DLC.
During the inference phase, all LoRA plugins are activated to extract their respective feature representations. These representations are then concatenated to form an integrated feature vector, which enhances the base model through a multi-task feature ensemble. For a linear classifier $\boldsymbol{W}$, the classification process can be formulated as:
$\hat{y}=\boldsymbol{W}[\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),...,\phi^{L_{T}}(\boldsymbol{x})]$ (9)
where $T$ denotes the maximum index of known tasks. Compared to the model expansion, LoRA plugin expansion can bring significant parameter advantages. Specifically, for a convolutional layer ($C_{in}$, $C_{out}$, $K\times K$), the space complexity of its parameters is $\Theta(C_{in}C_{out}K^{2})$, while a ConvLoRA only occupies $\Theta(r(C_{in}K^{2}+C_{out}))$. For expanding one layer of CNN over $T$ tasks, considering that in a CNN in CIL, $C_{out}$ is generally greater than or equal to $C_{in}$, we set $C_{in}=C_{out}=C$, and $K$ can be regarded as a constant (usually 3 or 7). Since $r$ is low-rank, it can also be regarded as a constant (generally set to 8 or 16). Then, using LoRA expansion reduces the parameter space complexity of expanding one layer of CNN from $\Theta(TC^{2})$ to $\Theta(TC)$. Considering that $C$ is generally between 64 and 512, reducing its order of magnitude from quadratic to linear can bring a huge reduction in parameters.
4.3 Dynamic Weighting Unit
For a sample from task $t$, the LoRA plugins trained on other tasks is considered non-target plugins. Given that the replay buffer contains only a minimal fraction of the original data for each task, these non-target plugins are not trained on the full data from task $t$. In particular, plugins trained on tasks 1 to t1t-1 never encounter samples from the task $t$. Consequently, the residuals extracted by these plugins could be largely irrelevant to the task $t$, thus introducing uninformative features or even detrimental noise to the classifier.
To mitigate the influence of uninformative residuals from non-target plugins, we will create a weighting unit corresponding to concentrated representations when training on a new task. This unit comprises two linear layers followed by a sigmoid function $\sigma(\cdot)$, denoted as $\omega(\mathbf{x})=\sigma\Big(\mathbf{W}_{2}\text{ReLU}\big(\mathbf{W}_{1}\mathbf{x}\big)\Big),\mathbf{x}\in\mathbb{R}^{k},\mathbf{W}_{1}\in\mathbb{R}^{k\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times k}$, where the sigmoid function normalizes the output to the [0, 1] range, indicating their importance weights. The concatenated representations are passed through this unit before being fed into the classifier, then weighted representations is derived from $\tilde{\mathbf{h}}^{(i)}=\mathbf{\omega}^{(i)}\odot\mathbf{h}^{(i)}$, where $\odot$ is element-wise multiplication operator. The weighting unit can be regarded as a gate for the classifier. Jointly trained with the classifier, it helps the model adapt to the high-dimensional feature space and automatically assigns lower weights to uninformative or noisy features.
The automatically trained weighting unit fails to explicitly model the varying importance of representations from different tasks. For a sample $i$ in task $t$, we denote the representations adapted by LoRA plugins trained on tasks 1 to t1t-1 as $\mathbf{h}^{(i)}_{\text{pre}}$ and those adapted by the remaining LoRA plugins as $\mathbf{h}^{(i)}_{\text{pos}}$, respectively. Then $\mathbf{h}^{(i)}_{\text{pre}}$ should be assigned lower weights as corresponding LoRA plugins were not optimized on task $t$. However, since the ground-truth task index of each sample is unavailable during the CIL testing, we cannot explicitly control the weighting unit based on task identity. Instead, we employ a supervised regression training to guide the weighting unit to approximate the desired weighting scheme. We decompose the weight $\mathbf{\omega}^{(i)}$ for sample $i$ into $\mathbf{\omega}^{(i)}_{\text{pre}},\mathbf{\omega}^{(i)}_{\text{pos}}$] Then the decision process of the linear classifier can be rewritten as follows:
$\boldsymbol{q}^{i}=\mathbf{W}_{\text{pre}}(\mathbf{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})+\mathbf{W}_{\text{pos}}(\mathbf{\omega}^{(i)}_{\text{pos}}\odot\mathbf{h}^{(i)}_{\text{pos}})$. (10)
Formally, the contribution of $\mathbf{h}^{(i)}_{\text{pre}}$ to the classifier’s logits can be expressed as $\mathbf{\xi}=\mathbf{W}_{\text{pre}}(\mathbf{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})$. To generate effective weights, an importance-aware optimization objective could be incorporated into the training process, denoted as $\mathcal{L}_{\text{IA}}$. This loss function employs the Mean Squared Error (MSE), expressed as follows:
$\mathcal{L}_{\text{IA}}=\frac{1}{k}{\|\mathbf{\omega}-\mathbf{\omega}_{\text{ideal}}\|}_{2},$ (11)
where the $\mathbf{\omega}_{\text{ideal}}$ is the ideal weight output, it has $\mathbf{\omega}^{(i)}_{\text{ideal-pos}}=1$ and $\mathbf{\omega}^{(i)}_{\text{ideal-pre}}=0$. In the ideal training case, the model has $\mathbf{\omega}^{(i)}_{\text{pos}}\to 1$ and $\mathbf{\omega}^{(i)}_{\text{pre}}\to 0$, then it has $\mathbb{E}[\mathbf{\xi}]\to 0$ and $\text{Var}[\mathbf{\xi}]\to 0$, so the $\mathbf{h}^{(i)}_{\text{pre}}$ have negligible influence on prediction logits. Even if the ideal scenario is hard to achieve, the classifier can still produce more accurate predictions as long as the representations adapted by reliable plugins are assigned higher weights, while those adapted by non-target plugins are appropriately suppressed. After enabling the weighting unit, $\mathcal{L}_{\text{IA}}$ is incorporated into the LoRA training, and $\mathcal{L}_{\text{LoRA}}$ in Eq. (8) is extended as follows:
$\mathcal{L}_{\text{LoRA}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}+\mathcal{L}_{\text{IA}}.$ (12)
Table 1: Last performance $A_{T}$ and average performance $\overline{A}$ comparison. ‘-DLC’ indicates the method using the DLC framework. ‘LoRA #P’ denotes the total parameter count of the LoRA plugins, and ‘Ext #P’ refers to the total parameter count of the feature extractor.
Methods CIFAR-100 Tiny-ImageNet ImageNet-100
B5 Inc5 B10 Inc10 B50 Inc10 B40 Inc40 B50 Inc10
$\overline{A}$ $_{T} A¯\overline{A}$ $_{T} A¯\overline{A}$ $_{T} A¯\overline{A}$ $_{T} A¯\overline{A}$ $A_{T}$
LWF 34.04 14.03 43.10 23.09 38.83 22.56 45.23 32.32 44.84 30.22
LWF-DLC 34.20 15.29 43.38 23.30 39.20 22.95 47.86 33.98 46.60 32.10
Replay 58.28 38.20 56.82 39.31 51.08 40.88 51.46 37.79 55.22 43.16
Replay-DLC 58.76 38.69 59.24 41.81 53.72 43.61 51.65 38.66 59.76 49.62
iCaRL 59.40 39.94 59.67 41.29 55.94 44.44 52.32 36.22 62.72 53.55
iCaRL-DLC 59.54 41.63 60.06 44.31 56.61 46.35 53.28 39.20 65.03 54.06
BiC 56.45 32.77 61.14 42.58 53.45 36.04 54.12 39.68 66.02 49.82
BiC-DLC 58.27 41.07 61.87 43.65 54.46 46.23 54.86 43.09 71.53 57.24
WA 62.72 45.96 65.82 51.80 65.04 56.94 53.68 39.69 65.53 56.55
WA-DLC 62.84 46.22 66.79 52.70 65.43 57.71 55.31 43.33 73.84 64.40
LoRA #P 102400 51200 30720 409600 491520
Ext #P 464154 464154 464154 11168832 11168832
5 Experiments
This section conducts extensive experiments. Sec. 5.2 evaluates the validity of DLC on replay and distillation-based baselines. Sec. 5.3 evaluates the compatibility of DLC on other data augmentation and plug-and-play loss function for CIL. Sec. 5.4 investigates the performance of baselines under memory-aligned case. An ablation study is conducted in 5.5 to analyze the contribution of the weighting unit.
5.1 Experimental Setup
Datasets and Split. We chose three benchmark datasets commonly used in CIL, including CIFAR-100 [krizhevsky2009learning], ImageNet-100 [deng2009imagenet], and Tiny-ImageNet-200 [vinyals2016matching], where the number following the dataset is the total number of classes included. Following the ‘B/Base-m, Inc-n’ rule proposed by [de2021continual], we slice each of the above three datasets into CIFAR-100 B10 Inc10, CIFAR-100 B50 Inc10, CIFAR-100 B5 INC5, ImageNet-100 B50 Inc10, and Tiny-ImageNet B40 Inc40 for the experiment, where $m$ is the number of classes in the first incremental task, and $n$ represents that of every remaining stage. For a fair comparison, we ensure that the training and test datasets are identical for all methods. Detailed information about the dataset is placed in the supplementary materials.
Implementation Details. All methods are implemented in PyTorch, with the baseline methods referencing the PyCIL [zhou2023pycil] and Gao et al. [gao2025maintaining]. We employ the lightweight ResNet-32 tailored for CIFAR datasets on CIFAR-100, while utilizing ResNet-18 for all other benchmark datasets. The optimizer employs SGD with a fixed learning rate of 0.1, incorporating a multi-step schedule. The batch size is configured at 128 for CIFAR-100 and 256 for all other datasets. The number of training epochs was set consistent with the baseline method. All the methods are run using the same random number seed.
Evaluation Metrics. Following the benchmark protocol settings [rebuffi2017icarl], we use $A_{b}$ to denote the $b$-stage accuracy on the test set that includes all known classes after trained with $D_{1},D_{2},\cdots,D_{b}$, $\overline{A}=\frac{1}{T}\sum_{b=1}^{T}A_{b}$ is average accuracy over $T$ tasks, and $A_{T}$ is the last accuracy on the test set that includes all learned tasks. We use $\overline{A}$ and $A_{T}$ to evaluate the model’s performance, which reflects the model’s dynamic learning capacity and generalization ability, respectively.
5.2 Evaluation on Non-Expansion Baselines
In this section, we evaluate the performance improvement brought by the proposed DLC framework when integrated with established distillation-based baselines. We select four representative methods—LwF [li2017learning], iCaRL [rebuffi2017icarl], WA [zhao2020maintaining], and BiC [wu2019large] —where LwF (without a replay buffer) is used to assess DLC’s enhancement effect in a pure distillation setting. Additionally, a simple Replay method is included to examine DLC’s contribution in scenarios relying solely on the herding rehearsal without distillation. We report the $\overline{A}$ and $A_{T}$ of these established methods with and without DLC across three datasets under varying task configurations. In all experiments, the replay buffer size is fixed at 2,000 exemplars. DLC expands only one LoRA plugin per task, applied to the last convolutional layer of the feature extractor. We also report the total parameters required by the LoRA plugins and the size of the entire feature extractor for comparison. The results are summarized in Table 1.
The results show that DLC significantly improves all classical replay or distillation-based methods with only one LoRA plugin per task. Moreover, the parameter overhead introduced by DLC is negligible. DLC achieves up to 8% accuracy improvement while expanding only 4.4% to 22% parameters equivalent to the feature extractor, highlighting its practical value in CIL. Fig. 3 shows the confusion matrix heatmap of iCaRL and iCaRL-DLC as an example. DLC increases the thermal intensity along the diagonal for most classes, demonstrating that it could enhance the model’s accuracy across most tasks.
Figure 3: Confusion matrix heatmaps of iCaRL with and w/o DLC on CIFAR-100 B10 Inc10.
5.3 Compatibility with other Plug-and-Play Enhancements
Figure 4: Comparison of $A_{T}$ with/without DLC on CIFAR-100 when enhanced with AutoAugment (AA) and $\mathcal{L}_{\text{inter-intra}}$.
In addition to DLC, other plug-and-play strategies have been proposed to enhance CIL methods. For example, the data augmentation technique AutoAugment [cubuk2018autoaugment] has been employed in the CIFAR-100 dataset by [wang2022foster, zheng2025task]. Meanwhile, two additional training losses $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ specifically designed for logit-based distillation methods have been introduced in [gao2025maintaining]. This subsection evaluates the compatibility of DLC with such enhance strategies. Specifically, we compare the accuracy of methods using AutoAugment and the loss $\mathcal{L}_{\text{inter-intra}}$ with and without DLC. As shown in the Fig. 4, DLC consistently brings significant improvements, demonstrating its compatibility with other plug-and-play enhancements.
5.4 Comparison under Fixed Memory Budget
Prior research [zhou2024class] has demonstrated that model-level expansion-based methods achieve state-of-the-art performance across all benchmarks. However, these approaches overlook the substantial memory overhead induced by dynamic architecture growth. Such methods implicitly introduce additional memory budgets, i.e., a model buffer for retaining parameters of the expanded model, which grants them a considerable and unaccounted advantage over other methods that do not store expansion models or adopt more lightweight expansion strategies.
Following the protocol established in [zhou2022model], this section presents a memory-aligned comparison of CIL baselines without pre-training. Under a fixed 23.5 MB memory budget, the memory not consumed by a method’s parameters is allocated to replay exemplars. To balance this, methods with fewer parameters store more exemplars, and vice versa. This ensures an equitable evaluation. In addition to the baselines used in Section 5.2, we add several state-of-the-art approaches for comprehensive comparison, including FOSTER [wang2022foster], DER [yan2021dynamically], MEMO [zhou2022model], and TagFex [zheng2025task]. Among them, DER, MEMO, and TagFex have incorporated the weight alignment from WA in their implementations. As summarized in Table 2, deploying DLC to the distillation-based method WA under the memory-aligned setting yields the highest classification accuracy.
Table 2: Performance comparison on CIFAR-100 with aligned memory cost. ‘#P’ represents the number of parameters (million). $\varepsilon$’ denotes the number of exemplars, and ‘MS’ denotes the memory size (MB). All methods use ResNet-32 as the backbone.
Method CIFAR100 B10 Inc10
MS #P $\varepsilon$ $\overline{A}$ $A_{T}$
iCaRL 23.5 0.46 7431 70.74 58.12
WA 23.5 0.46 7431 69.25 59.00
BiC 23.5 0.46 7431 70.29 59.20
FOSTER 23.5 0.46 7431 72.18 59.29
DER+WA 23.5 4.60 2000 71.24 60.26
MEMO+WA 23.5 3.62 3312 72.32 61.92
TagFex+WA 23.5 5.62 700 74.00 64.10
WA-DLC 23.5 0.72 7000 75.02 64.79
5.5 Ablation Study on the Weighting Unit
This subsection evaluates the impact of the proposed weighting unit on model performance. We compare the final accuracy of DLC-enhanced iCaRL, WA, and BiC with and without the weighting unit, as shown in Fig. 5. Results indicate that the proposed weighting unit contributes a certain accuracy gain. Although this improvement is relatively modest compared to the gains achieved by the LoRA plugin, the weighting unit could effectively direct the model’s attention to representations adapted by the pivotal LoRA plugin or key features of the sample, thereby enhancing their discriminative power. Specifically, we visualize the activation patterns of the last convolutional layer with the target task LoRA plugin using Grad-CAM [selvaraju2017grad], as illustrated in Fig. 6. For a “road” sample, despite both models correctly classifying it, the model equipped with the weighting unit effectively focuses on the road surface. Regarding a “pear” sample, when the weighting unit is ablated, the activation patterns remain inactive, indicating that representations adapted by the target LoRA contribute negligibly to classification. In contrast, with the weighting unit, the model not only correctly utilizes residuals from the target plugin but also focuses on discriminative regions, such as the cross‑section of the pear to improve the accuracy.
Figure 5: Comparison of $A_{T}$ with and without the weighting unit for DLC-enhanced methods on CIFAR-100.
Figure 6: Grad-CAM visualizations for the last convolutional layer integrated with the target-task LoRA plugin. The model is iCaRL-DLC trained on the CIFAR-100 B10 Inc10. ‘Original’ denotes the input image, while ‘W/O weight’ and ‘Weighting’ represent the results without and with the weighting unit.
6 Conclusion
While many class incremental learning methods avoid parameter-intensive expansion by relying on replay or distillation, they often face inherent performance limitations due to imbalanced replay buffers or constraints imposed by distillation. Our work presents DLC, a plug-and-play expansion method that enhances the performance of the base model by task-specific LoRA plugins. These plugins are trained separately and frozen to prevent interference from distillation loss and protect learned knowledge. A trainable weighting unit further suppresses irrelevant features from non-target plugins. Experimental results show that DLC achieves consistent enhancements in non-expansion CIL methods with a few parameter overhead. Future improvements could involve designing a more effective weighting unit to better integrate features or deploying LoRA in earlier layers for enhanced adaptability.

### v2

@@
-Parameter-Efficient Augment Plugin for Class-Incremental Learning
+Pushing the Limits of Distillation-Based Class-Incremental Learning via Lightweight Plugins
 Abstract
-Existing class-incremental learning (CIL) approaches based on replay or knowledge distillation are often constrained by forgetting or the stability-plasticity dilemma.
-Some expansion-based approaches could achieve higher accuracy.
-However, they always require significant parameter increases.
-In this paper, we propose a plugin extension paradigm termed the Deployment of extra LoRA Components (DLC) for non-pre-trained CIL scenarios.
-We treat the feature extractor trained through replay or distillation as a base model with rich knowledge.
+Existing replay and distillation-based class-incremental learning (CIL) methods are effective at retaining past knowledge but are still constrained by the stability–plasticity dilemma.
+Since their resulting models are learned over a sequence of incremental tasks, they encode rich representations and can be regarded as pre-trained bases.
+Building on this view, we propose a plug-in extension paradigm termed Deployment of LoRA Components (DLC) to enhance them.
 For each task, we use Low-Rank Adaptation (LoRA) to inject task-specific residuals into the base model’s deep layers.
@@
 This unit learns to assign importance scores to different LoRA-tuned representations.
-Like downloadable contents in software, our method serves as a plug-and-play enhancement that efficiently extends the base methods.
+Like downloadable content in software, DLC serves as a plug-and-play enhancement that efficiently extends the base methods.
 Remarkably, on the large-scale ImageNet‑100, with merely 4% of the parameters of a standard ResNet-18, our DLC model achieves a significant 8% improvement in accuracy, demonstrating exceptional efficiency.
-Moreover, it could surpass state-of-the-art methods under the fixed memory budget.
+Under a fixed memory budget, methods equipped with DLC surpass state-of-the-art expansion-based methods.
 1 Introduction
-In recent years, deep learning has achieved remarkable success across a variety of domains [jaimes2007multimodal, ye2019learning, chen2021large, chen2022learning, yang2024entaugment].
+In recent years, deep learning has achieved remarkable success across a variety of domains [14, 36, 5, 4, 34].
 However, most models rely on a key assumption that the data distribution and learned classes remain fixed.
-In real-world scenarios, data distributions evolve continuously and new classes emerge over time [gomes2017survey], making static training paradigms impractical.
-Moreover, learning each new task typically relies on introducing additional task-specific sub-models, leading to inefficient model growth over time.
-Class-Incremental Learning (CIL) [zhou2024class, liangloss, yang2025supervised] thus plays a crucial role in enabling models to continually learn from non-stationary data streams while retaining prior knowledge.
-Existing CIL methods  [zhou2024class] can be broadly categorized into rehearsal-based, restriction-based, and expansion-based approaches.
-Rehearsal-based methods maintain a small replay buffer by storing a limited number of representative exemplars from previous tasks [lopez2017gradient, rolnick2019experience, de2021rep].
-However, the number of stored exemplars is typically far smaller than that of samples in the current task, creating a severe data imbalance that biases learning toward new classes and thus exacerbates catastrophic forgetting.
-Restriction-based methods [nguyen2017variational, ahn2019uncertainty] aim to mitigate catastrophic forgetting by constraining model updates during new task learning.
-A representative example is knowledge distillation [rebuffi2017icarl, simon2021learning], which preserves previous knowledge by aligning the current model’s output with those of its earlier versions.
-Such constraints introduce a trade-off: the model must balance adherence to past knowledge with adaptation to new data.
-This stability-plasticity dilemma fundamentally limits the achievable performance of restriction-based approaches.
+In real-world scenarios, data distributions evolve continuously and new classes emerge over time [11], making static training paradigms impractical.
+Moreover, learning each new task typically relies on additional task-specific sub-models, leading to inefficient model growth over time.
+Class-Incremental Learning (CIL) [42, 17, 35] thus plays a crucial role in enabling models to continually learn from non-stationary data streams while retaining prior knowledge.
+Existing CIL methods  [42] can be broadly categorized into rehearsal-based, restriction-based, and expansion-based approaches.
+Rehearsal-based methods maintain a small replay buffer by storing a limited number of representative exemplars from previous tasks [18, 24, 8].
+Other categories of methods likewise adopt rehearsal mechanisms to retain exemplars from past tasks, thereby further improving model accuracy.
+Expansion-based CIL methods [19, 25, 32] allocate task-specific architectural blocks, creating isolated parameter spaces for new knowledge while freezing previously learned parameters.
+This strategy alleviates catastrophic forgetting and preserves plasticity on new tasks.
+However, rather than updating a single model incrementally, these methods expand the network and store, per task, an additional full [33] or partial (about 70%) [43] feature extractor, incurring substantial storage costs.
+For instance, with ResNet-18, each new task adds about 11.7MB of parameters, roughly equivalent to storing 3,998 images, which is about twice the replay buffer capacity under standard CIL settings.
+This overhead can limit their practicality for long task sequences, resource-constrained devices, and large-scale deployments.
 Figure 1: Parameter-accuracy comparison of different CIL methods on ImageNet-100 with ResNet-18.
@@
 iCaRL-DLC, WA-DLC, and BiC-DLC mean integrating DLC into the corresponding methods, respectively.
-To address this, expansion-based CIL methods [mallya2018piggyback, serra2018overcoming, xu2018reinforced] allocate task-specific architectural blocks, creating isolated parameter spaces for new knowledge and freezing previously learned parameters.
-While this isolates interference between tasks, the scalability limits as each new task requires its own feature extractor, causing parameters to grow linearly and incurring heavy storage costs.
-Recent pre-trained models-based methods [zhou2024continual] attempt to reduce this cost through lightweight fine-tuning modules.
-Yet, most PTM-based designs rely heavily on vision transformers with massive parameters (86M), and their fine-tuning modules can even surpass the size of conventional models.
-For instance, the introduction of a single Adapter module [zhou2024revisiting] incurs an parameter overhead (15M) greater than that of a complete ResNet-18 (11M).
-In addition, these PTM-based expansion methods face practical limitations, as PTMs are not universally available in practical applications.
-Therefore, expansion-based methods raise concerns despite strong performance.
-Their long-term sustainability and scalability in lifelong learning are questionable.
-This exposes a new quandary to CIL: How can we maintain high accuracy without a prohibitive linear increase in parameters as the number of tasks grows?
-To address this challenge, we propose the Deployment of extra LoRA Components (DLC).
-Rather than expanding the feature extractor model, DLC adopts a plug-and-play extension paradigm.
-Specifically, we attach task-specific Low-Rank Adaptation (LoRA) plugins to the deep layers of the base model.
-These plugins are updated exclusively during their corresponding task’s training phase, which occurs after the base model has been trained.
-This design ensures that the DLC plugins are trained independently, thus enabling their seamless and non-interfering integration into other methods.
-During inference, the base model loads all LoRA plugins to extract their feature representations, and they are aggregated into a composite one for final classification.
-However, non-target LoRA plugins that were not adequately updated on the dataset corresponding to the sample’s task may introduce noise or less informative residuals into the composite representation.
-So we propose a dedicated lightweight weighting unit before the classifier.
-This unit automatically assigns importance weights to different LoRA plugin outputs.
-DLC could enhance many non-expansion CIL approaches, such as knowledge distillation or replay methods.
-Experimental results on various benchmark datasets show that our method consistently improves the performance of replay- and distillation-based baselines.
-For instance, it could achieve up to 8% accuracy gains on ImageNet-100 using ResNet-18 and on CIFAR-100 using ResNet-32, while only increasing the total number of parameters by 4% and 11%, respectively.
-Furthermore, under the same memory budget, DLC outperforms existing state-of-the-art baselines.
+Restriction-based methods [21, 2] aim to mitigate catastrophic forgetting by constraining model updates on the new task.
+A representative example is knowledge distillation [23, 26], which preserves previous knowledge by aligning the current model’s output with its earlier versions.
+These methods perform CIL on a single model, leading to substantially lower storage overhead.
+As a result, these approaches are practically attractive in real-world settings.
+However, they inevitably introduce a trade-off [7]: the model must balance adherence to past knowledge with adaptation to new data.
+This stability-plasticity dilemma often becomes the key bottleneck for distillation-based approaches.
+Many enhancements [31, 38, 10] have been proposed to improve their accuracy, yet the core issue persists: distillation and new-task fitting are coupled objectives whose gradients can conflict, leaving performance governed by the same trade-off.
+A natural way to reduce this interference is to expand extra parameters for new tasks.
+However, scaling up the feature extractor is quickly parameter-prohibitive, and typical expansion-based schemes retrain task-specific extractors, breaking the shared single-model predictor that distillation relies on.
+This exposes a new quandary to CIL: How can we break the stability–plasticity dilemma by introducing a small set of task-specific auxiliary parameters, while remaining friendly to distillation-based methods?
+To address this challenge, we propose the Deployment of LoRA Components (DLC), a plug-and-play extension paradigm for CIL.
+We revisit parameter-efficient tuning modules in transfer learning beyond the frozen feature extractor assumption, and study them under the training dynamics of replay and distillation-based CIL.
+We attach task-specific lightweight plugins to deep layers of the base model and instantiate each plugin with LoRA in this work.
+These plugins provide additional knowledge by injecting residuals into the extracted features.
+For each incoming task, we first follow the underlying baseline to train the backbone while keeping all plugins frozen, and then update only the plugin associated with the current task.
+This decoupled update strategy trains each plugin independently, enabling DLC to be deployed into the baseline without interference.
+During inference, the model activates all plugins to extract their feature representations, which are aggregated for final classification.
+However, plugins that are not aligned with the input sample may contribute noisy or weakly informative residuals.
+We introduce a lightweight weighting unit before the classifier, which automatically assigns importance weights to different plugin outputs.
+Experimental results on multiple benchmarks show that DLC consistently improves replay and distillation-based baselines.
+For example, DLC achieves up to 8% accuracy gains on ImageNet-100 with ResNet-18 and on CIFAR-100 with ResNet-32, while increasing the total parameter count by only 4% to 30%.
+Under the same memory budget, DLC also outperforms existing state-of-the-art CIL baselines.
 Our contributions are summarized as follows:
-  • We propose a plugin extension framework DLC, which can be seamlessly integrated into diverse CIL methods without architectural modification by injecting task-specific residuals into deep layers via LoRA.
-  • We introduce a lightweight weighting unit for inference, which adaptively suppresses irrelevant LoRA activations according to their learned importance.
-This adaptive filtering mechanism effectively mitigates feature interference between old and new tasks, leading to more stable and accurate incremental classification.
+  • We propose DLC, a plug-and-play plugin extension framework that enables lightweight task-specific plugin residual injection in continually updated backbones, anchored by the consistent knowledge preserved through replay and distillation.
+  • We introduce a lightweight weighting unit that adaptively suppresses irrelevant plugin activations based on learned importance, reducing feature interference across tasks and leading to more stable and accurate incremental classification.
   • Extensive experiments show that DLC significantly boosts various CIL methods’ performance.
-Meanwhile, under the fixed memory budget, the DLC-enhanced distillation method could surpass SOTA baselines.
+Meanwhile, under the fixed memory budget, the DLC-enhanced distillation method could surpass SOTA CIL methods.
 2 Related Work
-DLC can be viewed as a specialized form of expansion-based CIL that augments base methods with task-specific plugins.
-This section first introduces methods facing the stability-plasticity dilemma that are compatible with DLC, followed by a review of expansion approaches.
-2.1 CIL with stability-plasticity dilemma
-To mitigate catastrophic forgetting, many methods implement constrained parameter updates during model optimization [nguyen2017variational, bian2024make, cui2025few].
-A prominent strategy in this category is knowledge distillation [simon2021learning, huang2024etag].
-It protects the previously acquired knowledge through a modified distillation loss from the old to the new model.
-iCaRL [rebuffi2017icarl] mitigates forgetting by distilling the logits from the old model into the new model.
-Subsequent methods like BiC [wu2019large] and WA [zhao2020maintaining] further propose bias correction and weight alignment techniques to calibrate the classifier.
-Moreover, these methods preserve extra knowledge of prior classes by employing replay buffers [shaul2024teal, aghasanli2025prototype] to save exemplars in training.
-Furthermore, recent research has developed plug-and-play auxiliary loss functions [gao2025maintaining] that can be readily integrated into existing methods to enhance their accuracy.
-However, the fixed capacity of the feature extractor poses a fundamental constraint.
-The model is inherently limited in its ability to retain knowledge from past tasks and learn new tasks effectively, forcing it to seek a balance between stability and plasticity through a trade-off between classification and distillation/regularization losses, which ultimately caps its achievable accuracy.
-2.2 Expansion-Based CIL
-Such methods expand specialized parameters for model [mallya2018piggyback, serra2018overcoming, xu2018reinforced].
-DER and TagFex [yan2021dynamically, zheng2025task] train an independent representation model on the new task for feature expansion, while the models of old tasks are frozen to retain the learned knowledge.
-MEMO [zhou2022model] splits the feature extractor into a generic shallow module and several task-specific deep modules.
-By increasing task-specific modules with fewer parameters, MEMO could save more memory space for storing replay samples than model-level expansion methods.
-Expansion-based methods achieve notably higher accuracy ceilings compared to fixed-capacity models.
-However, these methods result in a linear increase in the total number of parameters with tasks, imposing a significant storage overhead.
-Specifically, in the case of TagFex, the introduced attention module further increases the parameter overhead of the expansion.
-Even MEMO still expands over 70% of the extractor’s parameters for every new task.
-Emerging research leverages pre-trained models (PTMs) [zhou2024revisiting, mcdonnell2024ranpac, tan2024semantically] for CIL, mitigating catastrophic forgetting by leveraging knowledge encapsulated in PTMs.
-This has spurred the development of specialized parameter expansion methods designed for PTMs, which involve training lightweight fine-tuning modules like adapters [zhou2024expandable, sun2025mos].
-However, such methods paradoxically result in higher overall parameter counts rather than a reduction.
-This is primarily due to the large parameter size of ViT backbones—significantly exceeding that of ResNet—combined with the non-negligible overhead of their tuning modules.
-An adapter introduces parameters equivalent to 15% of the base ViT’s total.
-Furthermore, PTM is not available in many CIL cases.
-These parameter-efficient module expansion methods depend on the rich prior knowledge embedded in PTMs and cannot be directly transferred to non-pre-trained models.
+2.1 Class Incremental Learning
+Class-incremental learning primarily aims to prevent the model from forgetting old classes.
+Existing methods tackle this challenge from multiple perspectives, including rehearsal [16, 1] that replays representative exemplars from previous classes, regularization [3] or distillation [12] that limit overly aggressive updates on new data, expansion-based designs that train and retain task-specific full [33] or partial [43] feature extractors, and further integrate contrastive objectives [39] to stabilize representations.
+In addition, some methods leverage transfer learning with Transformers pretrained on large-scale datasets to provide stronger representations [40, 27].
+These methods often exhibit an accuracy–memory trade-off: approaches with higher recognition accuracy typically require storing more replay samples or allocating and maintaining additional model parameters through expansion or larger backbones.
+2.2 Distillation-Based CIL
+Distillation-based CIL trains the model on new tasks while regularizing it to match the predictions of the previous version.
+A wide range of distillation-based approaches have been explored, including combining distillation with replay [23], incorporating bias correction [31] or weight alignment [38], momentum-based distillation [20], and plug-and-play auxiliary objectives [10].
+Distillation remains appealing as it introduces minimal overhead, can leverage unlabeled incoming samples, and fits a simple single-model training loop, making it broadly applicable in practical settings such as online [30], imbalanced [22], and few-shot [37] CIL.
+However, the feature extractor’s fixed capacity imposes a fundamental constraint: the model must balance stability and plasticity by trading off classification and distillation losses, ultimately capping achievable accuracy.
 Figure 2: The proposed DLC framework.
@@
 3.1 Problem Formulation of CIL
-A CIL model aims to learn new classes while retaining previous knowledge from the continuously evolving data streams $D_{1},D_{2},\cdots,D_{t}$ [de2021continual], where $D_{i}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{n_{i}}$ containing $n_{i}$ instances.
+A CIL model aims to learn new classes while retaining previous knowledge from the continuously evolving data streams $D_{1},D_{2},\cdots,D_{t}$ [7], where $D_{i}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{n_{i}}$ containing $n_{i}$ instances.
 One class will appear only once on a given $D_{i}$ and will not be repeated.
-At each incremental learning stage $t$, only the current data stream $D_{t}$ and a few exemplars from the previous classes [rebuffi2017icarl] can be used for model training.
+At each incremental learning stage $t$, only the current data stream $D_{t}$ and a few exemplars from the previous classes [23] can be used for model training.
 For any stage $t$, the model has learned $D=D_{1}\cup D_{2}\cup\cdots\cup D_{t}$ streams, We aim to find a model $f(\mathbf{x}):\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes empirical risk across all learned classes, where $\mathcal{X}$ and $\mathcal{Y}=Y_{1}\cup Y_{2}\cdots\cup Y_{t}$ represent the sample’s feature and class spaces, respectively.
@@
 Concurrently, knowledge distillation functions as a functional regularizer, ensuring that the mappings for old classes remain stable by aligning the output logits (or distributions) with those of the prior model.
-Consequently, a model trained with the replay buffer and distillation retains significant knowledge about past classes, functioning as a capable but suboptimal ”generalist.” We posit that the knowledge stored in the model can be efficiently activated and tuned with minor task-specific parameters.
+Consequently, a model trained with replay and distillation preserves substantial knowledge of past classes, resulting in a single unified model that is broadly effective but suboptimal.
+We posit that the knowledge stored in the model can be efficiently activated with minor task-specific parameters.
 4 The Proposed Method
-While the use of rehearsal and restriction enable the model to retain knowledge from previous tasks, an additional objective or imbalanced exemplars could create an optimization trade-off.
-This results in a suboptimal convergence point for the cross-entropy loss or past data, ultimately capping the model’s classification performance.
-DLC proposes a distinctive paradigm, as shown in Fig.
+DLC proposes a plugin-level residual injection paradigm, as illustrated in Fig.
 2.
-We regard the model trained with a replay buffer and distillation as a foundation model, which stored knowledge across all tasks (albeit with suboptimal accuracy).
-Then refine it through residuals from each task-specific LoRA plugin to enhance accuracy.
-The details of the proposed DLC framework are elaborated in the following subsections.
-4.1 LoRA for Convolutional Layers
-A basic CIL model typically comprises a feature extractor $\phi(\boldsymbol{x})$ followed by a classifier $\boldsymbol{W}$.
-Most model expansion methods focus on expanding the $\phi(\boldsymbol{x})$ to achieve feature expansion.
-To maximize parameter efficiency, our approach avoids expanding the entire $\phi(\boldsymbol{x})$ or its deeper layers.
-Instead, we append parameter-efficient plugins after selected network layers within the $\phi(\boldsymbol{x})$ to adapt the representations produced by the base model.
-We specifically adopt Low-Rank Adaptation (LoRA) [hu2022lora] due to its minimal parameter size and its architectural flexibility.
-The standard LoRA method was originally designed for linear layers.
-Given that most non-pre-trained CIL models are based on Convolutional Neural Networks (CNNs), it is necessary to change the original LoRA formulation to the convolutional operations fundamental to CNNs.
-For a convolutional layer with weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times K\times K}$ and input $\boldsymbol{x}\in\mathbb{R}^{N\times C_{\text{in}}\times H\times W}$, the original forward pass is $\boldsymbol{h}=W_{\text{conv}}\ast\boldsymbol{x}$, where $\ast$ denotes the convolution operation.
-By analogy to LoRA, ConvLoRA constructs a low-rank adaptation for a convolutional weight tensor via a sequence of two convolutional operations.
-The forward pass first processes the input with $\mathbf{A}$:
-$\boldsymbol{z}=\mathbf{A}\ast\boldsymbol{x},$ (5)
-where $\mathbf{A}\in\mathbb{R}^{r\times C_{\text{in}}\times K\times K}$ projects the feature to a low-rank space. Then followed by kernel B:
-$\boldsymbol{h}_{\text{lora}}=\mathbf{B}\ast\boldsymbol{z},$ (6)
-where $B\in\mathbb{R}^{C_{\text{out}}\times r\times 1\times 1}$ restores channel dimensionality. In summary, the forward pass with ConvLoRA is:
-$\boldsymbol{h}^{\prime}=W_{\text{conv}}\ast\boldsymbol{x}+\frac{\alpha}{r}(B\ast(A\ast\boldsymbol{x})),$ (7)
-where $\alpha$ is a scaling hyperparameter, $r$ is the low-rank.
-4.2 Deploy LoRA as Task-Specific Plugins
-LoRA cannot be directly applied in non-PTM settings, not only because PTMs provide diverse knowledge.
-Without a frozen PTM backbone, the output of the model changes with parameter updates, leading to unstable layer inputs for LoRA.
-However, LoRA becomes applicable in replay- and distillation-based methods, as these approaches effectively constrain the variation of the model’s outputs on old tasks, thereby ensuring that the input features to LoRA remain close to the distribution encountered during its training phase.
-Specifically, after training on task $t$, the output deviation of the model at any layer $\ell$ will not exceed $K_{\ell}\Gamma_{t}$ (The detailed proof is provided in the supplementary material).
-Here, $K_{\ell}$ is determined by hyperparameters such as the network structure and distillation temperature, while $\Gamma_{t}$ characterizes the intensity of distribution discrepancy induced by selected distillation and replay strategies.
-This upper bound limits the maximum feature drift magnitude of the model on the old tasks’ distribution.
-Similar to expansion-based methods, our approach incorporates independent LoRA plugins for each task.
-Specifically, for each task $t$, we create a dedicated set of $k$ LoRA plugins for $\phi(\boldsymbol{x})$, denoted collectively as $L_{t}$.
+We treat a replay and distillation-trained network as a base model that accumulates knowledge across tasks but remains suboptimal in accuracy, and then augment it with task-specific plugins that inject residual knowledge into the extracted features to improve performance.
+In this work, each plugin is instantiated with a LoRA-style low-rank parameterization, but is used as a lightweight task extension rather than a conventional fine-tuning update.
+The proposed DLC is detailed in the following subsections.
+4.1 Task-Specific Plugins Deployment
+Parameter-efficient tuning is typically used in transfer learning by adapting a feature extractor with frozen parameters, where layer inputs remain largely stationary.
+Here, we revisit it beyond this static-backbone assumption and study it under replay- and distillation-based class-incremental learning, where old-task outputs are constrained, and intermediate feature drift is therefore controlled during continual updates.
+Concretely, after training on task $t$, the output deviation at layer $\ell$ is bounded by $K_{\ell}\Gamma_{t}$ (see Appendix), where $K_{\ell}$ depends on hyperparameters such as network structure and distillation temperature, and $\Gamma_{t}$ reflects the distribution discrepancy induced by the chosen replay and distillation strategy.
+This bound directly limits the worst-case feature drift on the old-task distribution, enabling stable optimization of lightweight auxiliary parameters in this dynamic regime.
+We introduce a plugin extension design and maintain an independent plugin set for each task.
+Specifically, for each task $t$, we create a dedicated set of $k$ plugins for $\phi(\boldsymbol{x})$, denoted as $L_{t}$.
 The $\phi(\boldsymbol{x})$ equipped with $L_{t}$ is denoted by $\phi^{L_{t}}(\boldsymbol{x})$.
-In DLC, LoRA plugins are deployed in the final layers of $\phi(\boldsymbol{x})$.
-Taking a CNN-based architecture as an example, the first LoRA plugin in $L_{t}$ is applied to the last convolutional layer of the $\phi(\boldsymbol{x})$.
-The remaining LoRA plugins are then sequentially deployed in a backward manner toward earlier layers.
-This design is motivated by two key considerations.
-First, features extracted from deeper layers are more semantically meaningful and are directly utilized for final classification decisions.
-Injecting task-specific residual adaptations via LoRA at this level allows for more direct enhancement of inter-class discriminability.
-Second, it demonstrates better compatibility with distillation-based objectives.
-When applied to shallow layers, the residual modifications introduced by LoRA would cause alterations in low-level features that propagate through the entire network, consequently amplifying the divergence in the final output logits.
-This increased discrepancy substantially raises the optimization difficulty for distillation losses, which rely on logit alignment between old and new models.
-In DLC, the LoRA plugins and $\phi(\boldsymbol{x})$ undergo a decoupled, two-phase training procedure.
-Specifically, upon the arrival of a new task $t$, we first train only the $\phi(\boldsymbol{x})$ and classifier $\boldsymbol{W}$, keeping all LoRA plugins frozen, to facilitate adaptation to the new task distribution.
-After this phase, $\phi(\boldsymbol{x})$ is frozen to prevent knowledge drift, and the $L_{t}$ are trained with the same data.
-Upon completion, $L_{t}$ is permanently frozen to safeguard the acquired task-specific knowledge from forgetting.
-The training objective for all LoRA plugins incorporates the cross-entropy loss and the auxiliary loss [yan2021dynamically, zhou2022model] commonly employed in expansion-based methods, formulated as follows:
-$\mathcal{L}_{\text{LoRA}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}.$ (8)
-The two-phase training procedure ensures that the LoRA plugins are trained without interference from distillation losses, enabling specialized learning for the current task and thereby improving its accuracy.
-Since the $\phi(\boldsymbol{x})$ remain frozen during LoRA updates, the training process of LoRA plugins exhibits stability.
-Furthermore, it prevents the $\phi(\boldsymbol{x})$’s training from being affected by LoRA, thereby maintaining the plug-and-play characteristic of DLC.
-During the inference phase, all LoRA plugins are activated to extract their respective feature representations.
+In DLC, plugins are deployed in the final layers of $\phi(\boldsymbol{x})$.
+Taking a CNN-based architecture as an example, the first plugin in $L_{t}$ is attached to the last convolutional layer of $\phi(\boldsymbol{x})$, and the remaining plugins are placed on progressively earlier layers moving backward through the network.
+This design is motivated by two considerations.
+First, deeper features are more semantically meaningful and are directly used for classification, so injecting task-specific residuals at this level can more effectively enhance inter-class discriminability.
+Second, it is more compatible with distillation-based objectives.
+If plugins are deployed in shallow layers, residual changes in low-level features would propagate through the network and amplify logit deviations, increasing the optimization difficulty of distillation losses that rely on output alignment between old and new models.
+In DLC, the plugins and $\phi(\boldsymbol{x})$ undergo a decoupled two-phase training procedure.
+Specifically, when a new task $t$ arrives, we first train the base model following the underlying baseline procedure, updating only $\phi(\boldsymbol{x})$ and the classifier $\boldsymbol{W}$ while keeping all plugins frozen.
+After this phase, $\phi(\boldsymbol{x})$ is frozen to prevent knowledge drift, and $L_{t}$ is trained with the same data.
+Upon completion, $L_{t}$ is permanently frozen to safeguard the acquired task-specific knowledge.
+The training objective for all plugins incorporates the cross-entropy loss and an auxiliary loss [33] commonly used in expansion-style methods, formulated as:
+$\mathcal{L}_{\text{plg}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}.$ (5)
+The two-phase procedure ensures that plugins are trained without interference from distillation losses, enabling specialized learning for the target task.
+Since $\phi(\boldsymbol{x})$ remains frozen during plugin updates, plugin training is stable, and it also prevents the baseline training of $\phi(\boldsymbol{x})$ from being affected by plugin training, preserving the plug-and-play property of DLC.
+During inference, all plugins are activated to extract their respective feature representations.
 These representations are then concatenated to form an integrated feature vector, which enhances the base model through a multi-task feature ensemble.
@@
 where $T$ denotes the maximum index of known tasks.
-Compared to the model expansion, LoRA plugin expansion can bring significant parameter advantages.
-Specifically, for a convolutional layer ($C_{in}$, $C_{out}$, $K\times K$), the space complexity of its parameters is $\Theta(C_{in}C_{out}K^{2})$, while a ConvLoRA only occupies $\Theta(r(C_{in}K^{2}+C_{out}))$.
+Notably, we do not need to run the full extractor $T$ times: each plugin injects a residual only at its deployment layer, so the backbone activations up to the earliest deployed layer are shared and computed once, avoiding redundant computation and improving inference efficiency.
+4.2 LoRA-Style Residual Plugin
+To instantiate each task-specific plugin with minimal parameters, we adopt a LoRA [13]-style low-rank residual parameterization.
+This design is well-suited for plugin deployment: the rank $r$ provides a simple capacity knob under a fixed budget, the same low-rank residual form applies to both linear and convolutional operators, and the residual structure supports our decoupled training and inference aggregation since plugins can be independently trained, frozen, and combined without modifying the backbone.
+We emphasize that the backbone weights are kept intact throughout, and the low-rank component is introduced as an auxiliary module that adds a task-specific residual to the feature.
+That is, the LoRA plugin does not replace or factorize the original weights, but injects additional task knowledge via an additive low-rank branch, which aligns with our plug-and-play deployment setting.
+Since most CIL pipelines train CNN-based backbones from scratch, we instantiate the plugin for convolutional operators using a LoRA-style low-rank residual branch (implemented as two convolutions).
+For a convolutional layer with weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times K\times K}$ and input $\boldsymbol{x}\in\mathbb{R}^{N\times C_{\text{in}}\times H\times W}$, the original forward pass is $\boldsymbol{h}=W_{\text{conv}}\ast\boldsymbol{x}$, where $\ast$ denotes the convolution operation.
+The LoRA plugin first processes the input with $\mathbf{A}$:
+$\boldsymbol{z}=\mathbf{A}\ast\boldsymbol{x},$ (7)
+where $\mathbf{A}\in\mathbb{R}^{r\times C_{\text{in}}\times K\times K}$ projects the feature to a low-rank space. This is followed by $\mathbf{B}$:
+$\boldsymbol{h}_{\text{lora}}=\mathbf{B}\ast\boldsymbol{z},$ (8)
+where $\mathbf{B}$ $\in\mathbb{R}^{C{\text{out}}\times r\times 1\times 1}$ restores the channel dimensionality, and $\boldsymbol{h}_{\text{lora}}$ is added as a residual to the layer output.
+In summary, the forward pass of a convolutional layer equipped with the LoRA-style residual plugin is:
+$\boldsymbol{h}^{\prime}=W_{\text{conv}}\ast\boldsymbol{x}+\frac{\alpha}{r}(\mathbf{B}\ast(\mathbf{A}\ast\boldsymbol{x})),$ (9)
+where $\alpha$ is a scaling hyperparameter, and $r$ is the rank.
+Compared to the model expansion, instantiating task-specific plugins with a LoRA-style parameterization can bring significant parameter advantages.
+Specifically, for a convolutional layer ($C_{in}$, $C_{out}$, $K\times K$), the space complexity of its parameters is $\Theta(C_{in}C_{out}K^{2})$, while the plugin only occupies $\Theta(r(C_{in}K^{2}+C_{out}))$.
 For expanding one layer of CNN over $T$ tasks, considering that in a CNN in CIL, $C_{out}$ is generally greater than or equal to $C_{in}$, we set $C_{in}=C_{out}=C$, and $K$ can be regarded as a constant (usually 3 or 7).
 Since $r$ is low-rank, it can also be regarded as a constant (generally set to 8 or 16).
-Then, using LoRA expansion reduces the parameter space complexity of expanding one layer of CNN from $\Theta(TC^{2})$ to $\Theta(TC)$.
-Considering that $C$ is generally between 64 and 512, reducing its order of magnitude from quadratic to linear can bring a huge reduction in parameters.
+Then, using LoRA-style plugin instantiation reduces the parameter-space complexity of expanding a single CNN layer from $\Theta(TC^{2})$ to $\Theta(TC)$.
+Considering that $C$ is generally between 64 and 512, reducing its order of magnitude from quadratic to linear can substantially reduce the number of parameters.
 4.3 Dynamic Weighting Unit
-For a sample from task $t$, the LoRA plugins trained on other tasks is considered non-target plugins.
+For a sample from task $t$, all the plugins trained on other tasks are considered non-target plugins.
 Given that the replay buffer contains only a minimal fraction of the original data for each task, these non-target plugins are not trained on the full data from task $t$.
@@
 Even if the ideal scenario is hard to achieve, the classifier can still produce more accurate predictions as long as the representations adapted by reliable plugins are assigned higher weights, while those adapted by non-target plugins are appropriately suppressed.
-After enabling the weighting unit, $\mathcal{L}_{\text{IA}}$ is incorporated into the LoRA training, and $\mathcal{L}_{\text{LoRA}}$ in Eq.
+After enabling the weighting unit, $\mathcal{L}_{\text{IA}}$ is incorporated into the plugin training, and $\mathcal{L}_{\text{plg}}$ in Eq.
 (5) is extended as follows:
-$\mathcal{L}_{\text{LoRA}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}+\mathcal{L}_{\text{IA}}.$ (12)
+$\mathcal{L}_{\text{plg}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}+\mathcal{L}_{\text{IA}}.$ (12)
 Table 1: Last performance $A_{T}$ and average performance $\overline{A}$ comparison.
 ‘-DLC’ indicates the method using the DLC framework.
-‘LoRA #P’ denotes the total parameter count of the LoRA plugins, and ‘Ext #P’ refers to the total parameter count of the feature extractor.
+‘Extra/Feat #P’ denotes the total number of extra parameters required for deploying DLC and the feature extractor, respectively.
 Methods CIFAR-100 Tiny-ImageNet ImageNet-100
@@
 $\overline{A}$ $_{T} A¯\overline{A}$ $_{T} A¯\overline{A}$ $_{T} A¯\overline{A}$ $_{T} A¯\overline{A}$ $A_{T}$
-LWF 34.04 14.03 43.10 23.09 38.83 22.56 45.23 32.32 44.84 30.22
-LWF-DLC 34.20 15.29 43.38 23.30 39.20 22.95 47.86 33.98 46.60 32.10
 Replay 58.28 38.20 56.82 39.31 51.08 40.88 51.46 37.79 55.22 43.16
@@
 WA-DLC 62.84 46.22 66.79 52.70 65.43 57.71 55.31 43.33 73.84 64.40
-LoRA #P 102400 51200 30720 409600 491520
-Ext #P 464154 464154 464154 11168832 11168832
+MKD 56.86 36.87 55.73 36.83 46.68 36.17 49.34 34.64 65.64 54.68
+MKD-DLC 57.51 37.33 58.10 40.40 52.18 40.99 50.29 35.75 66.42 56.23
+DAKD 60.01 40.15 60.69 42.32 54.33 43.09 51.98 36.73 60.76 51.50
+DAKD-DLC 60.47 41.54 61.01 43.81 56.99 46.00 52.78 37.38 64.73 53.94
+Extra/Feat #P (MB) 0.15 0.46 0.07 0.46 0.04 0.46 0.45 11.17 0.52 11.17
 5 Experiments
@@
 Sec.
-5.2 evaluates the validity of DLC on replay and distillation-based baselines.
+5.2 evaluates the validity of DLC on distillation-based baselines.
 Sec.
-5.3 evaluates the compatibility of DLC on other data augmentation and plug-and-play loss function for CIL.
+5.3 investigates the performance of SOTA CIL methods under the memory-aligned case.
 Sec.
-5.4 investigates the performance of baselines under memory-aligned case.
-An ablation study is conducted in 5.5 to analyze the contribution of the weighting unit.
+5.4 evaluates the compatibility of DLC with other data augmentation and plug-and-play loss functions for CIL.
+Sec.
+5.5 analyzes the additional inference overhead introduced by DLC in terms of FLOPs and latency.
 5.1 Experimental Setup
 Datasets and Split.
-We chose three benchmark datasets commonly used in CIL, including CIFAR-100 [krizhevsky2009learning], ImageNet-100 [deng2009imagenet], and Tiny-ImageNet-200 [vinyals2016matching], where the number following the dataset is the total number of classes included.
-Following the ‘B/Base-m, Inc-n’ rule proposed by [de2021continual], we slice each of the above three datasets into CIFAR-100 B10 Inc10, CIFAR-100 B50 Inc10, CIFAR-100 B5 INC5, ImageNet-100 B50 Inc10, and Tiny-ImageNet B40 Inc40 for the experiment, where $m$ is the number of classes in the first incremental task, and $n$ represents that of every remaining stage.
+We chose three benchmark datasets commonly used in CIL, including CIFAR-100 [15], ImageNet-100 [9], and Tiny-ImageNet-200 [28], where the number following the dataset is the total number of classes included.
+Following the ‘B/Base-m, Inc-n’ rule proposed by [7], we slice each of the above three datasets into CIFAR-100 B10 Inc10, CIFAR-100 B50 Inc10, CIFAR-100 B5 Inc5, ImageNet-100 B50 Inc10, and Tiny-ImageNet B40 Inc40 for the experiment, where $m$ is the number of classes in the first incremental task, and $n$ represents that of every remaining stage.
 For a fair comparison, we ensure that the training and test datasets are identical for all methods.
-Detailed information about the dataset is placed in the supplementary materials.
+Detailed information about the dataset is placed in the Appendix section.
 Implementation Details.
-All methods are implemented in PyTorch, with the baseline methods referencing the PyCIL [zhou2023pycil] and Gao et al.
-[gao2025maintaining].
+All methods are implemented in PyTorch, with the baseline methods referencing the PyCIL [41], and Gao et al.
+[10].
 We employ the lightweight ResNet-32 tailored for CIFAR datasets on CIFAR-100, while utilizing ResNet-18 for all other benchmark datasets.
 The optimizer employs SGD with a fixed learning rate of 0.1, incorporating a multi-step schedule.
-The batch size is configured at 128 for CIFAR-100 and 256 for all other datasets.
-The number of training epochs was set consistent with the baseline method.
+The detailed training hyperparameters, like the optimizer and batch size, largely follow the default settings in PyCIL.
 All the methods are run using the same random number seed.
+To eliminate potential confounding effects from increased training epochs, all baselines are run under both the PyCIL-recommended epoch setting and the total epoch budget of DLC’s two-stage training, and the maximum accuracy is reported.
 Evaluation Metrics.
-Following the benchmark protocol settings [rebuffi2017icarl], we use $A_{b}$ to denote the $b$-stage accuracy on the test set that includes all known classes after trained with $D_{1},D_{2},\cdots,D_{b}$, $\overline{A}=\frac{1}{T}\sum_{b=1}^{T}A_{b}$ is average accuracy over $T$ tasks, and $A_{T}$ is the last accuracy on the test set that includes all learned tasks.
+Following the benchmark protocol settings [23], we use $A_{b}$ to denote the $b$-stage accuracy on the test set that includes all known classes after trained with $D_{1},D_{2},\cdots,D_{b}$, $\overline{A}=\frac{1}{T}\sum_{b=1}^{T}A_{b}$ is average accuracy over $T$ tasks, and $A_{T}$ is the last accuracy on the test set that includes all learned tasks.
 We use $\overline{A}$ and $A_{T}$ to evaluate the model’s performance, which reflects the model’s dynamic learning capacity and generalization ability, respectively.
-5.2 Evaluation on Non-Expansion Baselines
-In this section, we evaluate the performance improvement brought by the proposed DLC framework when integrated with established distillation-based baselines.
-We select four representative methods—LwF [li2017learning], iCaRL [rebuffi2017icarl], WA [zhao2020maintaining], and BiC [wu2019large] —where LwF (without a replay buffer) is used to assess DLC’s enhancement effect in a pure distillation setting.
-Additionally, a simple Replay method is included to examine DLC’s contribution in scenarios relying solely on the herding rehearsal without distillation.
+5.2 Effectiveness on Distillation-Based CIL
+In this section, we evaluate the performance improvement brought by the proposed DLC framework across different distillation-based CIL objectives, including iCaRL [23], WA [38], BiC [31], DAKD [12], and MKD [20].
+Additionally, a simple herding-based Replay method [23] is included to examine DLC’s contribution in scenarios relying solely on the rehearsal without distillation.
 We report the $\overline{A}$ and $A_{T}$ of these established methods with and without DLC across three datasets under varying task configurations.
@@
 DLC expands only one LoRA plugin per task, applied to the last convolutional layer of the feature extractor.
-We also report the total parameters required by the LoRA plugins and the size of the entire feature extractor for comparison.
+We also report the total parameters required by DLC and the size of the entire feature extractor for comparison.
 The results are summarized in Table 1.
@@
 Moreover, the parameter overhead introduced by DLC is negligible.
-DLC achieves up to 8% accuracy improvement while expanding only 4.4% to 22% parameters equivalent to the feature extractor, highlighting its practical value in CIL.
+DLC achieves up to 8% accuracy improvement while expanding only 4.8% to 32% parameters equivalent to the feature extractor, highlighting its practical value in CIL.
 Fig.
@@
 DLC increases the thermal intensity along the diagonal for most classes, demonstrating that it could enhance the model’s accuracy across most tasks.
+In addition, we compare the final accuracy of DLC-enhanced iCaRL, WA, and BiC with and without the weighting unit, as shown in Fig.
+4.
+Results indicate that the proposed weighting unit contributes a certain accuracy gain.
 Figure 3: Confusion matrix heatmaps of iCaRL with and w/o DLC on CIFAR-100 B10 Inc10.
-5.3 Compatibility with other Plug-and-Play Enhancements
-Figure 4: Comparison of $A_{T}$ with/without DLC on CIFAR-100 when enhanced with AutoAugment (AA) and $\mathcal{L}_{\text{inter-intra}}$.
-In addition to DLC, other plug-and-play strategies have been proposed to enhance CIL methods.
-For example, the data augmentation technique AutoAugment [cubuk2018autoaugment] has been employed in the CIFAR-100 dataset by [wang2022foster, zheng2025task].
-Meanwhile, two additional training losses $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ specifically designed for logit-based distillation methods have been introduced in [gao2025maintaining].
-This subsection evaluates the compatibility of DLC with such enhance strategies.
-Specifically, we compare the accuracy of methods using AutoAugment and the loss $\mathcal{L}_{\text{inter-intra}}$ with and without DLC.
-As shown in the Fig.
-4, DLC consistently brings significant improvements, demonstrating its compatibility with other plug-and-play enhancements.
+Figure 4: Comparison of $A_{T}$ with and without the weighting unit for DLC-enhanced methods on CIFAR-100.
 5.3 Comparison under Fixed Memory Budget
-Prior research [zhou2024class] has demonstrated that model-level expansion-based methods achieve state-of-the-art performance across all benchmarks.
-However, these approaches overlook the substantial memory overhead induced by dynamic architecture growth.
-Such methods implicitly introduce additional memory budgets, i.e., a model buffer for retaining parameters of the expanded model, which grants them a considerable and unaccounted advantage over other methods that do not store expansion models or adopt more lightweight expansion strategies.
-Following the protocol established in [zhou2022model], this section presents a memory-aligned comparison of CIL baselines without pre-training.
-Under a fixed 23.5 MB memory budget, the memory not consumed by a method’s parameters is allocated to replay exemplars.
-To balance this, methods with fewer parameters store more exemplars, and vice versa.
-This ensures an equitable evaluation.
-In addition to the baselines used in Section 5.2, we add several state-of-the-art approaches for comprehensive comparison, including FOSTER [wang2022foster], DER [yan2021dynamically], MEMO [zhou2022model], and TagFex [zheng2025task].
-Among them, DER, MEMO, and TagFex have incorporated the weight alignment from WA in their implementations.
-As summarized in Table 2, deploying DLC to the distillation-based method WA under the memory-aligned setting yields the highest classification accuracy.
 Table 2: Performance comparison on CIFAR-100 with aligned memory cost.
@@
 TagFex+WA 23.5 5.62 700 74.00 64.10
-WA-DLC 23.5 0.72 7000 75.02 64.79
-5.5 Ablation Study on the Weighting Unit
-This subsection evaluates the impact of the proposed weighting unit on model performance.
-We compare the final accuracy of DLC-enhanced iCaRL, WA, and BiC with and without the weighting unit, as shown in Fig.
-5.
-Results indicate that the proposed weighting unit contributes a certain accuracy gain.
-Although this improvement is relatively modest compared to the gains achieved by the LoRA plugin, the weighting unit could effectively direct the model’s attention to representations adapted by the pivotal LoRA plugin or key features of the sample, thereby enhancing their discriminative power.
-Specifically, we visualize the activation patterns of the last convolutional layer with the target task LoRA plugin using Grad-CAM [selvaraju2017grad], as illustrated in Fig.
+WA-DLC 23.5 0.54 7327 75.02 64.79
+Prior research [42] has demonstrated that model-level expansion-based methods achieve state-of-the-art performance across all standard CIL benchmarks.
+However, these approaches overlook the substantial memory overhead induced by dynamic architecture growth.
+Such methods implicitly introduce additional memory budgets, i.e., a model buffer for retaining parameters of the expanded model, which grants them a considerable and unaccounted advantage over other methods that do not store expansion models or adopt more lightweight expansion strategies.
+Following the protocol established in [43], this section presents a memory-aligned comparison of CIL baselines.
+Under a fixed 23.5 MB memory budget, the memory not consumed by a method’s parameters is allocated to replay exemplars.
+To balance this, methods with fewer parameters store more exemplars, and vice versa.
+This ensures an equitable evaluation.
+In addition to the baselines used in Section 5.2, we add several state-of-the-art approaches for comprehensive comparison, including FOSTER [29], DER [33], MEMO [43], and TagFex [39].
+Among them, DER, MEMO, and TagFex have incorporated the weight alignment from WA in their implementations.
+As summarized in Table 2, deploying DLC to the distillation-based method WA under the memory-aligned setting yields the highest classification accuracy.
+5.4 Compatibility with other Plug-and-Play Enhancements
+Figure 5: Comparison of $A_{T}$ with/without DLC on CIFAR-100 when enhanced with AutoAugment (AA) and $\mathcal{L}_{\text{inter-intra}}$.
+In addition to DLC, other plug-and-play strategies have been proposed to enhance CIL methods.
+For example, the data augmentation technique AutoAugment [6] has been employed in the CIFAR-100 dataset by [29, 39].
+Meanwhile, two additional training losses $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ specifically designed for logit-based distillation methods have been introduced in [10].
+This subsection evaluates the compatibility of DLC with such enhanced strategies.
+Specifically, we compare the accuracy of methods using AutoAugment and the loss $\mathcal{L}_{\text{inter-intra}}$ with and without DLC.
+For experiments with $\mathcal{L}_{\text{inter-intra}}$, we follow the original implementation and use ResNet-18 as the backbone.
+As shown in the Fig.
+5, DLC consistently brings significant improvements, demonstrating its compatibility with other plug-and-play enhancements.
+5.5 Additional Computational Overhead from DLC
+This section investigates the additional inference cost introduced by deploying DLC.
+Using iCaRL as a representative baseline, we evaluate two deployment variants: iCaRL equipped with one LoRA plugin per task, and iCaRL equipped with both the LoRA plugins and the weighting unit.
+Concretely, on ImageNet-100, we report the total FLOPs and the average inference latency per sample (ms) measured on the same test set.
+The results are shown in Fig.
 6.
-For a “road” sample, despite both models correctly classifying it, the model equipped with the weighting unit effectively focuses on the road surface.
-Regarding a “pear” sample, when the weighting unit is ablated, the activation patterns remain inactive, indicating that representations adapted by the target LoRA contribute negligibly to classification.
-In contrast, with the weighting unit, the model not only correctly utilizes residuals from the target plugin but also focuses on discriminative regions, such as the cross‑section of the pear to improve the accuracy.
-Figure 5: Comparison of $A_{T}$ with and without the weighting unit for DLC-enhanced methods on CIFAR-100.
-Figure 6: Grad-CAM visualizations for the last convolutional layer integrated with the target-task LoRA plugin.
-The model is iCaRL-DLC trained on the CIFAR-100 B10 Inc10.
-‘Original’ denotes the input image, while ‘W/O weight’ and ‘Weighting’ represent the results without and with the weighting unit.
+Overall, neither FLOPs nor latency increases substantially as the number of tasks grows.
+Instead, both grow mildly as the task sequence extends.
+Even under long sequences such as B5 Inc5, the increases in FLOPs and inference time are only about 4%.
+These results indicate that DLC provides a lightweight and computation-friendly parameter extension for enhancing distillation-based CIL.
+Figure 6: Comparison of inference FLOPs and per-sample inference latency with and w/o DLC under different task settings.
 6 Conclusion
-While many class incremental learning methods avoid parameter-intensive expansion by relying on replay or distillation, they often face inherent performance limitations due to imbalanced replay buffers or constraints imposed by distillation.
-Our work presents DLC, a plug-and-play expansion method that enhances the performance of the base model by task-specific LoRA plugins.
-These plugins are trained separately and frozen to prevent interference from distillation loss and protect learned knowledge.
-A trainable weighting unit further suppresses irrelevant features from non-target plugins.
-Experimental results show that DLC achieves consistent enhancements in non-expansion CIL methods with a few parameter overhead.
-Future improvements could involve designing a more effective weighting unit to better integrate features or deploying LoRA in earlier layers for enhanced adaptability.
+In this work, we address a challenge in distillation-based class-incremental learning: improving the model performance with only a small amount of task-specific auxiliary parameters, without breaking the single-model distillation setting.
+We propose DLC, a plug-and-play plugin extension framework that augments replay and distillation-based baselines with task-specific LoRA-style residual plugins.
+DLC trains each plugin in a decoupled manner and then freezes it, which both avoids interference from distillation objectives and preserves the acquired task knowledge.
+We further introduce a lightweight weighting unit to down-weight irrelevant plugin features at inference.
+Extensive experiments demonstrate that DLC consistently improves replay and distillation-based CIL baselines with few parameter overhead.
+Future work will explore more expressive weighting designs for stronger feature integration, as well as earlier-layer plugin deployment to further improve adaptability.
+Appendix
+Appendix A Dataset Description
+CIFAR-100 CIFAR-100 (Canadian Institute for Advanced Research 100-class dataset) is a widely used standard benchmark dataset in the field of computer vision, particularly for image classification tasks.
+It is an expanded version of the CIFAR-10 dataset, curated by the Canadian Institute for Advanced Research.
+A defining feature of this dataset is its fine-grained categorical structure.
+It contains a total of 100 distinct fine-grained classes (fine labels), organized hierarchically under 20 broader superclasses (coarse labels), with each superclass encompassing precisely 5 fine-grained classes.
+This hierarchical organization provides a valuable foundation for researching multi-level classification.
+The dataset comprises 60,000 tiny color images, each with a fixed resolution of 32x32 pixels.
+This compact size makes the dataset manageable for rapid experimentation and algorithm iteration, while simultaneously presenting significant challenges for model recognition capabilities due to the extreme low resolution.
+The 60,000 images are explicitly divided into 50,000 training images and 10,000 test images, meaning each fine-grained class includes 500 training images and 100 test images.
+CIFAR-100 offers broad and diverse category coverage, encompassing common objects (various vehicles, furniture, appliances), animals (mammals, fish, insects), plants (flowers, trees), and natural scenes (mountains, oceans), among others.
+Its extensive number of classes combined with its moderate scale make it an ideal testbed for evaluating and comparing the performance and generalization capabilities of image classification models—especially Convolutional Neural Networks (CNNs)—as well as for exploring advanced tasks like few-shot learning and transfer learning.
+ImageNet-100 ImageNet-100 is a commonly used subset of the well-known large-scale image dataset, ImageNet.
+Its purpose is to significantly reduce computational resource requirements and training time while preserving much of ImageNet’s richness and challenge, making it more suitable for rapid prototyping, algorithm exploration, and research in resource-constrained environments.
+As the name implies, ImageNet-100 consists of 100 classes (synsets) selected from the original ImageNet dataset.
+The selection of these classes is typically not entirely random; it often follows specific strategies (e.g., choosing semantically distinct classes or classes with relatively higher sample counts) to ensure the subset remains representative and diverse.
+The dataset contains approximately 100,000 high-resolution color images curated from these 100 classes.
+This represents a substantial reduction in scale compared to the millions of images in the full ImageNet.
+The standard partition of the dataset usually includes a training set and a validation set.
+The training set contains roughly 1,000 images per class, totaling approximately 100,000 images.
+The validation set contains about 50 images per class, totaling around 5,000 images.
+The images themselves are typically of variable sizes (significantly larger than CIFAR’s 32x32 pixels) and require preprocessing (like cropping and resizing) before being fed into models.
+ImageNet-100 inherits core characteristics from its parent dataset: rich and realistic image content, high semantic distinction between classes, significant intra-class variation (the same object under different viewpoints, lighting, occlusion), and inter-class similarities.
+It is widely used for training and evaluating image classification models, as well as for transfer learning tasks.
+Compared to smaller datasets like CIFAR, models pre-trained on ImageNet-100 generally learn more generalizable visual features suitable for fine-tuning on downstream tasks.
+Compared to the full ImageNet, it offers faster iteration cycles and lower hardware barriers to entry.
+TinyImageNet-200 TinyImageNet-200 is a downsized variant of the ImageNet dataset, specifically created for educational and experimental purposes in deep learning and computer vision research.
+It features a curated subset of 200 distinct object categories (synsets) carefully selected from the full ImageNet hierarchy.
+The dataset contains roughly 100,000 high-resolution color images – significantly fewer than its parent dataset but substantially more than smaller benchmarks like CIFAR.
+Each image in Tiny ImageNet has been consistently downscaled and cropped to a fixed resolution of 64x64 pixels, maintaining a usable size while greatly reducing computational demands compared to processing full-sized ImageNet images.
+The dataset comes with a predefined partition: a training set containing 500 images per class, a separate validation set with 50 images per class, and a test set also containing 50 images per class, totaling approximately 100,000 training images, 10,000 validation images, and 10,000 test images.
+Alongside the images, TinyImageNet provides essential metadata files including a wnids.txt file listing the WordNet IDs for all 200 classes, a words.txt file mapping these IDs to human-readable labels, and a structured val_annotations.txt file listing the ground-truth classes for validation images.
+Its manageable scale, balanced class representation, and standardized structure make Tiny ImageNet-200 a popular choice for academic courses exploring deep learning architectures (particularly CNNs and Vision Transformers), as well as for research prototyping on tasks such as image classification, transfer learning, and model robustness evaluation, especially where computational resources are constrained but larger challenge than CIFAR is needed.
+Appendix B Feature Drift Bounds under Logit Distillation and Herding Replay
+Before applying methods like LoRA to continually updated models, it is imperative to guarantee that the feature extractor’s output deviation for previous tasks remains bounded before and after training.
+Otherwise, if catastrophic forgetting causes a significant shift in outputs, the plugin modules would be unable to extract meaningful residuals.
+Setting.
+Consider a sequence of tasks $D_{1},D_{2},\dots,D_{t}$. At task $t$, the model is trained on
+$D_{\mathrm{train}}=D_{t}\cup D_{\exp},$ (13)
+where $D_{\exp}$ is the class-balanced exemplar set from previous tasks based on the herding replay strategy in iCaRL.
+Let $\theta$ denote the parameters of a convolutional feature extractor $\phi(\boldsymbol{x})$, and let $h_{\ell}(x;\theta)\in\mathbb{R}^{d_{\ell}}$ denote the output of layer $\ell$ in $\phi(\boldsymbol{x})$.
+We denote the parameters of $\phi(\boldsymbol{x})$ as $\theta_{t-1}$ and $\theta_{t}$ for before and after training on task $t$.
+We adopt the logit-based distillation approach as a representative case for analysis.
+To constrain forgetting, the logits are distilled from the teacher to the student.
+The logits of teacher and student are denoted as $\hat{z}(x)$ and $z(x)$.
+For temperature $T>0$, define the softened probabilities
+$\hat{q}^{T}(x)=\mathrm{softmax}\!\big(\hat{z}(x)/T\big),$ (14)
+$q^{T}(x)=\mathrm{softmax}\!\big(z(x)/T\big).$ (15)
+The cross-entropy-based distillation loss is
+$\mathcal{L}_{\mathrm{KD}}(x)=-\sum_{i=1}^{C}\hat{q}^{T}_{i}(x)\log q^{T}_{i}(x)$ (16)
+$=H\!\big(\hat{q}^{T}(x)\big)+\mathrm{KL}\!\big(\hat{q}^{T}(x)\|q^{T}(x)\big).$ (17)
+Where $H(\hat{q}^{T}(x))$ denotes the Shannon entropy of the teacher distribution,
+$H(\hat{q}^{T}(x)):=-\sum_{i=1}^{C}\hat{q}^{T}_{i}(x)\log\hat{q}^{T}_{i}(x),$
+so the KD loss decomposes into the teacher’s entropy and the KL divergence between the teacher and the student.
+The empirical distillation loss over $D_{\exp}$ is
+$\bar{L}_{\mathrm{KD}}=\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\mathcal{L}_{\mathrm{KD}}(x).$ (18)
+It serves as a measurable proxy for how closely the student imitates the teacher on the subset of old-task samples preserved through herding replay.
+Our goal is to study how much the intermediate-layer features drift after learning on a new task. For any layer $\ell$, define the drift as:
+$\Delta h_{\ell}(x)=h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1}).$ (19)
+which quantifies the amount by which the feature extractor changes its representation of an old-task sample.
+Bounding this is crucial for ensuring that LoRA plugins trained on earlier feature distributions remain effective after the model is updated on new tasks.
+In addition, to compare logits in a way that reflects changes in their meaningful structure rather than arbitrary additive shifts, we use centered logits:
+$\bar{z}(x)=z(x)-\tfrac{1}{C}\sum_{k=1}^{C}z_{k}(x)\mathbf{1},$ (20)
+$\bar{\hat{z}}(x)=\hat{z}(x)-\tfrac{1}{C}\sum_{k=1}^{C}\hat{z}_{k}(x)\mathbf{1}.$ (21)
+Centered logit is necessary because logits are defined only up to a constant offset: adding a constant to all logit entries leaves the softmax distribution unchanged.
+By removing this offset, we ensure that the norms of logit differences reflect genuine changes in class-relative predictions rather than uninformative bias shifts.
+This step allows the later Lipschitz bounds to connect logit discrepancies with feature discrepancies in a well-defined way.
+Assumptions
+A1 (Non-degenerate teacher distribution). There exists $\delta\in(0,1/C]$ that satisfies:
+$\hat{q}^{T}_{i}(x)\geq\delta\quad\text{for all }x\in D_{\exp}\text{ and }i.$ (22)
+This condition prevents extreme logit sharpness and ensures the log-softmax function is well-conditioned.
+The assumption is directly satisfied in standard knowledge distillation settings.
+In particular, the use of a temperature parameter $T>1$ produces softened teacher probabilities,
+$\hat{q}^{T}_{i}(x)=\frac{\exp(\hat{z}_{i}(x)/T)}{\sum_{j}\exp(\hat{z}_{j}(x)/T)},$
+which guarantees that no class probability collapses to zero.
+Alternatively, label smoothing also enforces a uniform lower bound on class probabilities.
+Thus, A1 follows from standard KD practice and is not a restrictive condition.
+A2 (Local bi-Lipschitz continuity).
+Let $\Phi_{\ell}$ denote the mapping from the $\ell$-th layer feature to the centered logits $\bar{z}$ through subsequent layers.
+There exist constants $0<\mu_{\ell}\leq L_{\ell}<\infty$ such that
+$\mu_{\ell}\|u-v\|_{2}\leq\|\Phi_{\ell}(u)-\Phi_{\ell}(v)\|_{2}\leq L_{\ell}\|u-v\|_{2},$ (23)
+for all $u,v$ within a neighborhood of past-task features in $D_{\exp}$.
+This ensures logits and features change proportionally, which is well supported by the structure of modern neural networks.
+Convolutional layers, affine transformations, and common nonlinearities (ReLU, SiLU, GELU) are all Lipschitz continuous.
+Since compositions of Lipschitz functions remain Lipschitz, the entire feature-to-logit map possesses an upper Lipschitz constant.
+The lower Lipschitz constant reflects local non-collapse of the representation: empirically, trained deep networks do not map distinct inputs from the same manifold region to identical logits.
+Similar assumptions have been used in analyses of embedding stability and invertibility of deep models.
+Therefore, A2 expresses a standard geometric regularity of neural networks rather than a special structural requirement.
+A3 (Empirical KD value).
+At the end of training, the empirical KD loss over $D_{\exp}$ equals $\bar{L}_{\mathrm{KD}}$.
+This is not an additional assumption but rather a definition of the observable optimization outcome.
+The analysis treats $\bar{L}_{\mathrm{KD}}$ as a known scalar summarizing how well the student matches the teacher on replay samples.
+A4 (Generalization from herded exemplars).
+The replay buffer $D_{\exp}$ maintains an equal number of samples per class, and $\mathcal{L}_{\mathrm{KD}}(x)\in[0,\log C]$ is bounded, so with probability at least $1-\delta_{0}$, we have:
+$\big|\mathbb{E}_{x\sim\mathcal{D}_{<t}}[\mathcal{L}_{\mathrm{KD}}(x)]-\bar{L}_{\mathrm{KD}}\big|\leq\epsilon_{\mathrm{gen}},$ (24)
+$\epsilon_{\mathrm{gen}}=O\!\Big(\sqrt{\tfrac{\log(1/\delta_{0})}{|D_{\exp}|}}\Big),$ (25)
+where $\mathcal{D}_{<t}$ is the joint distribution of all previous tasks.
+This assumption holds because the following properties guarantee that the empirical distillation loss concentrates around its population expectation.
+The boundedness of the distillation loss,
+$0\leq\mathcal{L}_{\mathrm{KD}}(x)\leq\log C,$
+implies that empirical averages of $\mathcal{L}_{\mathrm{KD}}$ satisfy standard concentration inequalities.
+Moreover, the herding procedure constructs $D_{\exp}$ as a class-balanced set that approximates class means in feature space, making it an approximately representative subset of the past-task distribution $\mathcal{D}_{<t}$.
+As a result, the empirical quantity
+$\bar{L}_{\mathrm{KD}}=\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\mathcal{L}_{\mathrm{KD}}(x)$
+forms a low-variance estimator of the population expectation. Applying Hoeffding-type bounds for bounded losses yields, with probability at least $1-\delta_{0}$,
+$\big|\mathbb{E}_{x\sim\mathcal{D}_{<t}}[\mathcal{L}_{\mathrm{KD}}(x)]-\bar{L}_{\mathrm{KD}}\big|\;\leq\;O\!\left(\sqrt{\tfrac{\log(1/\delta_{0})}{|D_{\exp}|}}\right),$
+Auxiliary Lemmas
+Lemma 1 (KD controls probability discrepancy). For any $x$, by Pinsker’s inequality:
+$\|q^{T}(x)-\hat{q}^{T}(x)\|_{1}\leq\sqrt{2\,\mathrm{KL}(\hat{q}^{T}(x)\|q^{T}(x))}\leq\sqrt{2(\mathcal{L}_{\mathrm{KD}}(x)-H(\hat{q}^{T}(x)))}.$ (26)
+Averaging this equation over $D_{\exp}$ gives
+$\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\|q^{T}(x)-\hat{q}^{T}(x)\|_{1}\leq\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\tfrac{1}{|D_{\exp}|}\sum_{x}H(\hat{q}^{T}(x))\Big)}.$
+Lemma 2 (Probability discrepancy controls logit discrepancy). Softmax relates logits and probabilities by:
+$(z_{i}-z_{j})=T\log(\frac{q^{T}_{i}}{q^{T}_{j}}).$ (27)
+So under A1, for any $i\neq j$, Because of Lipschitz continuity of $\log p$ over a bounded domain, differences in log-probabilities are linearly bounded by differences in probabilities:
+$|\log q^{T}_{i}-\log\hat{q}^{T}_{i}|\leq\frac{T}{\delta}(|q^{T}_{i}-\hat{q}^{T}_{i}|).$ (28)
+So it has:
+$|(z_{i}-z_{j})-(\hat{z}_{i}-\hat{z}_{j})|=T\Big|\log\frac{q^{T}_{i}}{q^{T}_{j}}-\log\frac{\hat{q}^{T}_{i}}{\hat{q}^{T}_{j}}\Big|$ (29)
+$\leq\frac{T}{\delta}(|q^{T}_{i}-\hat{q}^{T}_{i}|+|q^{T}_{j}-\hat{q}^{T}_{j}|).$ (30)
+Therefore,
+$\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\leq\frac{T}{\delta}\|q^{T}(x)-\hat{q}^{T}(x)\|_{1}.$ (31)
+Averaging Eq.(31) over $D_{\exp}$ and combining it with Eq. (B) in Lemma 1 yields:
+$\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\leq\frac{T}{\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\tfrac{1}{|D_{\exp}|}\sum_{x}H(\hat{q}^{T}(x))\Big)}.$ (32)
+Lemma 3. Under A2, for any $x$, it has:
+$\mu_{\ell}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}\leq\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}$ (33)
+$\leq L_{\ell}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}.$ (34)
+Main Conclusions
+Theorem 1 (Empirical bound on feature drift). For any layer $\ell$,
+$\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}\leq\frac{T}{\mu_{\ell}\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\tfrac{1}{|D_{\exp}|}\sum_{x}H(\hat{q}^{T}(x))\Big)}.$ (35)
+Proof. This conclusion can be straightforwardly proved by combining the Eq.(33) in Lemma 3 (first averaging it over $D_{\exp}$ ) with the Eq.(32) of Lemma 2.
+Theorem 2 (Distributional bound over past-task distribution). Under A1–A4, with probability at least $1-\delta_{0}$,
+$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}\leq\frac{T}{\mu_{\ell}\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\mathbb{E}_{x\sim\mathcal{D}_{<t}}[H(\hat{q}^{T}(x))]+\epsilon_{\mathrm{gen}}\Big)}.$ (36)
+Proof. The expectation form of Lemma 2 states that
+$\ \mathbb{E}_{x\sim\mathcal{D}_{<t}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\;\leq\;\frac{T}{\delta}\sqrt{2\Big(\mathbb{E}_{x\sim\mathcal{D}_{<t}}\mathcal{L}_{\mathrm{KD}}(x)-\mathbb{E}_{x\sim\mathcal{D}_{<t}}H(\hat{q}^{T}(x))\Big)}.$ (37)
+A4 ensures that the empirical distillation loss generalizes to the past-task distribution:
+$\Big|\mathbb{E}_{x\sim\mathcal{D}_{<t}}\mathcal{L}_{\mathrm{KD}}(x)-\bar{L}_{\mathrm{KD}}\Big|\;\leq\;\epsilon_{\mathrm{gen}}.$ (38)
+Hence,
+$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\mathcal{L}_{\mathrm{KD}}(x)\;\leq\;\bar{L}_{\mathrm{KD}}+\epsilon_{\mathrm{gen}}.$ (39)
+Substituting this into Eq.(37) yields:
+$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\;\leq\;\frac{T}{\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\mathbb{E}_{x\sim\mathcal{D}_{<t}}H(\hat{q}^{T}(x))+\epsilon_{\mathrm{gen}}\Big)}.$ (40)
+Lemma 3 states that for every $x$,
+$\mu_{\ell}\big\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\big\|_{2}\;\leq\;\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}.$ (41)
+Taking expectation on both sides gives:
+$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\big\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\big\|_{2}\;\leq\;\frac{1}{\mu_{\ell}}\mathbb{E}_{x\sim\mathcal{D}_{<t}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}.$ (42)
+Then the conclusion to be proved can be obtained by combining Eq.(40) with Eq.(42).
+Theorems 1–2 imply
+$\mathbb{E}\|\Delta h_{\ell}(x)\|\leq\underbrace{\tfrac{T}{\mu_{\ell}\delta}}_{\text{constant }K_{\ell}}\times\underbrace{\sqrt{2(\Delta_{\mathrm{KD}}+\epsilon_{\mathrm{gen}})}}_{\text{variable }\Gamma_{t}},$ (43)
+where $\Delta_{\mathrm{KD}}=\mathbb{E}[\mathcal{L}_{\mathrm{KD}}(x)-H(\hat{q}^{T}(x))]$ reflects the effectiveness of distillation on the data of previous tasks.
+A smaller $\Delta_{\mathrm{KD}}$ represents the predictive behavior of the student model that aligns with that of the teacher model on these tasks.
+$\epsilon_{\mathrm{gen}}=O(\sqrt{\log(1/\delta_{0})/|D_{\exp}|})$ decreases as more exemplars are stored.
+The constant $\tfrac{T}{\mu_{\ell}\delta}$ depends on network smoothness and temperature parameters.
+The $K_{\ell}$ captures the conditioning of the feature-to-logit map and the softness of the teacher distribution, while the $\Gamma_{t}$ measures how much logit drift the distillation and replay objective permits.
+This bound shows that when distillation is optimized and the exemplars approximate the old-task distribution well, the deep feature representations of past data remain stable even as the model learns new tasks.
+Therefore, even though $\phi(\boldsymbol{x})$ continuously updates its parameters, the output of each layer for previous tasks remains relatively stable.
+This allows the plugins trained on earlier tasks to stay effective, thereby enabling the deployment of DLC in replay- and distillation-based CIL methods.

### v3

@@
-Pushing the Limits of Distillation-Based Class-Incremental Learning via Lightweight Plugins
+Pushing the Limits of Distillation-Based Continual Learning via Classifier-Proximal Lightweight Plugins
 Abstract
-Existing replay and distillation-based class-incremental learning (CIL) methods are effective at retaining past knowledge but are still constrained by the stability–plasticity dilemma.
-Since their resulting models are learned over a sequence of incremental tasks, they encode rich representations and can be regarded as pre-trained bases.
-Building on this view, we propose a plug-in extension paradigm termed Deployment of LoRA Components (DLC) to enhance them.
-For each task, we use Low-Rank Adaptation (LoRA) to inject task-specific residuals into the base model’s deep layers.
-During inference, representations with task-specific residuals are aggregated to produce classification predictions.
-To mitigate interference from non-target LoRA plugins, we introduce a lightweight weighting unit.
-This unit learns to assign importance scores to different LoRA-tuned representations.
-Like downloadable content in software, DLC serves as a plug-and-play enhancement that efficiently extends the base methods.
-Remarkably, on the large-scale ImageNet‑100, with merely 4% of the parameters of a standard ResNet-18, our DLC model achieves a significant 8% improvement in accuracy, demonstrating exceptional efficiency.
-Under a fixed memory budget, methods equipped with DLC surpass state-of-the-art expansion-based methods.
+Continual learning requires models to learn continuously while preserving prior knowledge under evolving data streams.
+Distillation-based methods are appealing for retaining past knowledge in a shared single-model framework with low storage overhead.
+However, they remain constrained by the stability-plasticity dilemma: knowledge acquisition and preservation are still optimized through coupled objectives, and existing enhancement methods do not alter this underlying bottleneck.
+To address this issue, we propose a plugin extension paradigm termed Distillation-aware Lightweight Components (DLC) for distillation-based CL.
+DLC deploys lightweight residual plugins into the base feature extractor’s classifier-proximal layer, enabling semantic-level residual correction for better classification accuracy while minimizing disruption to the overall feature extraction process.
+During inference, plugin-enhanced representations are aggregated to produce classification predictions.
+To mitigate interference from non-target plugins, we further introduce a lightweight weighting unit that learns to assign importance scores to different plugin-enhanced representations.
+DLC could deliver a significant 8% accuracy gain on large-scale benchmarks while introducing only a 4% increase in backbone parameters, highlighting its exceptional efficiency.
+Moreover, DLC is compatible with other plug-and-play CL enhancements and delivers additional gains when combined with them.
 1 Introduction
-In recent years, deep learning has achieved remarkable success across a variety of domains [14, 36, 5, 4, 34].
-However, most models rely on a key assumption that the data distribution and learned classes remain fixed.
-In real-world scenarios, data distributions evolve continuously and new classes emerge over time [11], making static training paradigms impractical.
-Moreover, learning each new task typically relies on additional task-specific sub-models, leading to inefficient model growth over time.
-Class-Incremental Learning (CIL) [42, 17, 35] thus plays a crucial role in enabling models to continually learn from non-stationary data streams while retaining prior knowledge.
-Existing CIL methods  [42] can be broadly categorized into rehearsal-based, restriction-based, and expansion-based approaches.
-Rehearsal-based methods maintain a small replay buffer by storing a limited number of representative exemplars from previous tasks [18, 24, 8].
-Other categories of methods likewise adopt rehearsal mechanisms to retain exemplars from past tasks, thereby further improving model accuracy.
-Expansion-based CIL methods [19, 25, 32] allocate task-specific architectural blocks, creating isolated parameter spaces for new knowledge while freezing previously learned parameters.
-This strategy alleviates catastrophic forgetting and preserves plasticity on new tasks.
-However, rather than updating a single model incrementally, these methods expand the network and store, per task, an additional full [33] or partial (about 70%) [43] feature extractor, incurring substantial storage costs.
-For instance, with ResNet-18, each new task adds about 11.7MB of parameters, roughly equivalent to storing 3,998 images, which is about twice the replay buffer capacity under standard CIL settings.
-This overhead can limit their practicality for long task sequences, resource-constrained devices, and large-scale deployments.
-Figure 1: Parameter-accuracy comparison of different CIL methods on ImageNet-100 with ResNet-18.
-DLC achieves the best balance between parameter efficiency and accuracy compared to prior SOTA methods.
+In recent years, deep learning has achieved remarkable success across a broad spectrum of multimedia applications [39, 4, 10, 35, 41].
+However, most existing models are built on the assumption that data distributions and semantic categories remain fixed after training.
+In real-world multimedia scenarios, data streams evolve continuously, and novel categories emerge over time [9], making static training paradigms increasingly inadequate.
+Meanwhile, naively adapting models to newly arrived data often requires additional specific modules, leading to inefficient model growth and limited scalability.
+Continual learning (CL) [47, 20, 40] therefore plays a crucial role in enabling deep learning models to continuously adapt to non-stationary data while preserving prior knowledge.
+Existing CL methods [47] can be broadly categorized into rehearsal-based, restriction-based, and expansion-based approaches.
+Rehearsal-based methods maintain a small replay buffer by storing a limited number of representative exemplars from earlier learning stages [22, 28, 6].
+Other categories of methods likewise adopt rehearsal mechanisms to retain exemplars from previous learning stages, thereby further improving model accuracy.
+Expansion-based CL methods [23, 29, 37] allocate stage-specific architectural blocks, creating isolated parameter spaces for newly acquired knowledge while freezing previously learned parameters.
+This strategy alleviates catastrophic forgetting and preserves plasticity during subsequent learning stages.
+However, rather than updating a single model progressively, these methods expand the network and store, for each learning stage, an additional full [38] or partial (about 70%) [48] feature extractor, incurring substantial storage costs.
+For instance, with ResNet-18, each new learning stage adds about 11.7MB of parameters, roughly equivalent to storing 3,998 images, which is about twice the replay buffer capacity under standard CL settings.
+This overhead can limit their practicality for continual learning.
+Figure 1: Parameter-accuracy comparison of different CL methods on ImageNet-100.
+DLC achieves the best balance between parameter efficiency and accuracy compared to prior distillation-based and SOTA expansion-based methods.
 iCaRL-DLC, WA-DLC, and BiC-DLC mean integrating DLC into the corresponding methods, respectively.
-Restriction-based methods [21, 2] aim to mitigate catastrophic forgetting by constraining model updates on the new task.
+Restriction-based methods [25, 2] aim to mitigate catastrophic forgetting by constraining model updates during new learning stages.
 A representative example is knowledge distillation [27, 30], which preserves previous knowledge by aligning the current model’s output with its earlier versions.
-These methods perform CIL on a single model, leading to substantially lower storage overhead.
+These methods perform CL on a single model, leading to substantially lower storage overhead.
 As a result, these approaches are practically attractive in real-world settings.
-However, they inevitably introduce a trade-off [7]: the model must balance adherence to past knowledge with adaptation to new data.
+However, they inevitably introduce a trade-off [5]: the model must balance adherence to past knowledge with adaptation to new data, since the distillation and knowledge acquisition are coupled objectives whose gradients can conflict.
 This stability-plasticity dilemma often becomes the key bottleneck for distillation-based approaches.
-Many enhancements [31, 38, 10] have been proposed to improve their accuracy, yet the core issue persists: distillation and new-task fitting are coupled objectives whose gradients can conflict, leaving performance governed by the same trade-off.
-A natural way to reduce this interference is to expand extra parameters for new tasks.
-However, scaling up the feature extractor is quickly parameter-prohibitive, and typical expansion-based schemes retrain task-specific extractors, breaking the shared single-model predictor that distillation relies on.
-This exposes a new quandary to CIL: How can we break the stability–plasticity dilemma by introducing a small set of task-specific auxiliary parameters, while remaining friendly to distillation-based methods?
-To address this challenge, we propose the Deployment of LoRA Components (DLC), a plug-and-play extension paradigm for CIL.
-We revisit parameter-efficient tuning modules in transfer learning beyond the frozen feature extractor assumption, and study them under the training dynamics of replay and distillation-based CIL.
-We attach task-specific lightweight plugins to deep layers of the base model and instantiate each plugin with LoRA in this work.
-These plugins provide additional knowledge by injecting residuals into the extracted features.
-For each incoming task, we first follow the underlying baseline to train the backbone while keeping all plugins frozen, and then update only the plugin associated with the current task.
+Many enhancement methods [36, 43, 8] have been proposed to improve their accuracy, but most of them mainly introduce additional training objectives, leaving the core stability-plasticity dilemma fundamentally unresolved, leaving performance governed by the same trade-off.
+A natural way to reduce this interference is to expand extra parameters.
+However, scaling up the feature extractor is quickly parameter-prohibitive, and expanding multiple extractors breaks the shared single-model predictor that distillation relies on.
+This exposes a new quandary to CL: How can we break the stability–plasticity dilemma by introducing a small set of auxiliary parameters, while remaining friendly to distillation-based methods?
+To address this challenge, we propose the Distillation-aware Lightweight Components (DLC), a plug-and-play extension paradigm for CL.
+We revisit parameter-efficient tuning modules in transfer learning and study them under the training dynamics of distillation-based CL, where the backbone is not a knowledge-rich pretrained model but a dynamically evolved learner trained over continual stages.
+We attach several lightweight plugins to the classifier-proximal layer of the base feature extractor, i.e., its final feature layer before classification.
+These plugins perform specific semantic calibration by injecting residual corrections into classifier-proximal features.
+When adding each plugin, we first follow the underlying baseline to train the backbone while keeping all plugins frozen, and then freeze the backbone to train only the plugin introduced for the current stage.
 This decoupled update strategy trains each plugin independently, enabling DLC to be deployed into the baseline without interference.
@@
 For example, DLC achieves up to 8% accuracy gains on ImageNet-100 with ResNet-18 and on CIFAR-100 with ResNet-32, while increasing the total parameter count by only 4% to 30%.
-Under the same memory budget, DLC also outperforms existing state-of-the-art CIL baselines.
+Under the same memory budget, DLC also outperforms existing state-of-the-art CL baselines.
 Our contributions are summarized as follows:
-  • We propose DLC, a plug-and-play plugin extension framework that enables lightweight task-specific plugin residual injection in continually updated backbones, anchored by the consistent knowledge preserved through replay and distillation.
-  • We introduce a lightweight weighting unit that adaptively suppresses irrelevant plugin activations based on learned importance, reducing feature interference across tasks and leading to more stable and accurate incremental classification.
-  • Extensive experiments show that DLC significantly boosts various CIL methods’ performance.
-Meanwhile, under the fixed memory budget, the DLC-enhanced distillation method could surpass SOTA CIL methods.
+  • We propose DLC, a plug-and-play plugin extension framework that deploys lightweight residual plugins at classifier-proximal layers in continually updated backbones, enabling effective semantic correction while preserving compatibility with distillation-preserved knowledge.
+  • We introduce a lightweight weighting unit that adaptively suppresses irrelevant plugin activations based on learned importance, leading to more stable and accurate continual learning.
+  • Extensive experiments show that DLC significantly boosts the performance of various CL methods and remains compatible with other plug-and-play enhancement techniques.
+Meanwhile, under a fixed memory budget, the DLC-enhanced distillation method can surpass state-of-the-art expansion-based CL methods.
 2 Related Work
-2.1 Class Incremental Learning
-Class-incremental learning primarily aims to prevent the model from forgetting old classes.
-Existing methods tackle this challenge from multiple perspectives, including rehearsal [16, 1] that replays representative exemplars from previous classes, regularization [3] or distillation [12] that limit overly aggressive updates on new data, expansion-based designs that train and retain task-specific full [33] or partial [43] feature extractors, and further integrate contrastive objectives [39] to stabilize representations.
-In addition, some methods leverage transfer learning with Transformers pretrained on large-scale datasets to provide stronger representations [40, 27].
-These methods often exhibit an accuracy–memory trade-off: approaches with higher recognition accuracy typically require storing more replay samples or allocating and maintaining additional model parameters through expansion or larger backbones.
-2.2 Distillation-Based CIL
-Distillation-based CIL trains the model on new tasks while regularizing it to match the predictions of the previous version.
-A wide range of distillation-based approaches have been explored, including combining distillation with replay [23], incorporating bias correction [31] or weight alignment [38], momentum-based distillation [20], and plug-and-play auxiliary objectives [10].
-Distillation remains appealing as it introduces minimal overhead, can leverage unlabeled incoming samples, and fits a simple single-model training loop, making it broadly applicable in practical settings such as online [30], imbalanced [22], and few-shot [37] CIL.
-However, the feature extractor’s fixed capacity imposes a fundamental constraint: the model must balance stability and plasticity by trading off classification and distillation losses, ultimately capping achievable accuracy.
+2.1 Continual learning
+Continual learning primarily aims to prevent the model from forgetting acquired knowledge.
+Existing methods tackle this challenge from multiple perspectives.
+Rehearsal-based methods [19, 1] replays representative exemplars from previous classes.
+Regularization [3] or distillation-based methods [27, 36, 43, 13] limit overly aggressive updates on new data.
+Expansion-based methods train and retain task-specific full [38] or partial [48] feature extractors, and further integrate contrastive objectives [44] or pre-trained models [45, 31] for better representations.
+These methods often exhibit a performance–memory trade-off: approaches with higher overall performance typically require storing more replay samples or allocating and maintaining additional model parameters through expansion.
+2.2 Distillation-Based CL
+Distillation-based CL trains the model on new data while regularizing it to match the predictions of the previous version.
+A wide range of distillation-based approaches have been explored, including combining distillation with replay [27], incorporating bias correction [36] or weight alignment [43], momentum-based distillation [24], and plug-and-play enhancement methods [3, 8, 18].
+Distillation remains appealing as it introduces minimal overhead, can leverage unlabeled incoming samples, and fits a simple single-model training loop, making it broadly applicable in practical settings [34, 26, 42].
+However, the feature extractor’s fixed capacity imposes a fundamental constraint: the model must balance stability and plasticity by trading off new-data learning and distillation losses, ultimately capping achievable performance.
 Figure 2: The proposed DLC framework.
 Left: Training.
-When a new task $t$ arrives, a dedicated plugin set $L_{t}$ is created.
-Then sequentially train the feature extractor and $L_{t}$.
+When a new $D_{t}$ arrives, a dedicated plugin $L_{t}$ is created.
+First follows the base method to train the backbone, and then freezes the backbone to train $L_{t}$.
 Right: Test.
-The feature extractor sequentially loads all task plugins $L_{t}$ to produce enhanced representations, which are concatenated, weighted by a gating unit, and classified.
+Using all plugins to enhance the feature extractor representations, then reweight by a weighting unit, concatenated, and classified.
 3 Preliminaries
-3.1 Problem Formulation of CIL
-A CIL model aims to learn new classes while retaining previous knowledge from the continuously evolving data streams $D_{1},D_{2},\cdots,D_{t}$ [7], where $D_{i}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{n_{i}}$ containing $n_{i}$ instances.
-One class will appear only once on a given $D_{i}$ and will not be repeated.
-At each incremental learning stage $t$, only the current data stream $D_{t}$ and a few exemplars from the previous classes [23] can be used for model training.
-For any stage $t$, the model has learned $D=D_{1}\cup D_{2}\cup\cdots\cup D_{t}$ streams, We aim to find a model $f(\mathbf{x}):\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes empirical risk across all learned classes, where $\mathcal{X}$ and $\mathcal{Y}=Y_{1}\cup Y_{2}\cdots\cup Y_{t}$ represent the sample’s feature and class spaces, respectively.
+3.1 Problem Formulation
+A CL model aims to learn new data while retaining previous knowledge from the continuously evolving data streams $D_{1},D_{2},\cdots,D_{t}$, where $D_{i}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{n_{i}}$ containing $n_{i}$ instances.
+At each learning stage $t$, the current data stream $D_{t}$ is treated as a task, and training is performed using only $D_{t}$ with a few exemplars from the previous stage [27].
+After the model advances to the next learning stage, all earlier stages, except for the stored exemplars, will become inaccessible.
+When there are test samples for inference, the model needs to make predictions based on the label spaces of all the tasks it has encountered.
+We do not provide the task ID of the sample to the model, making this a more challenging class-incremental continual learning [5] setup.
+For any stage $t$, the model has learned $D=D_{1}\cup D_{2}\cup\cdots\cup D_{t}$ tasks, We aim to find a model $f(\mathbf{x}):\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes empirical risk across all learned data, where $\mathcal{X}$ and $\mathcal{Y}=Y_{1}\cup Y_{2}\cdots\cup Y_{t}$ represent the sample’s feature and label spaces, respectively.
 The effectiveness of the model is evaluated by the test dataset $D^{test}=D_{1}^{test}\cup D_{2}^{test}\cup\cdots\cup D_{t}^{test}$, which can be expressed in Eq.(1):
@@
 3.2 –
-Consistent Knowledge Preserved in CIL Considering the classic CIL scenario with only a single feature extractor.
-The objective function used in non-CIL settings is the cross-entropy loss $\mathcal{L}_{\text{CE}}$, while restriction-based CIL methods incorporate an additional restriction term $\mathcal{L}_{\text{MEM}}$ that enforces the preservation of knowledge acquired from previous tasks.
+Knowledge in Distillation-Based CL Considering the classic CL scenario with only a single feature extractor.
+The objective function used in non-CL settings is the cross-entropy loss $\mathcal{L}_{\text{CE}}$, while distillation-based CL methods incorporate an additional restriction term $\mathcal{L}_{\text{KD}}$ that enforces the preservation of knowledge acquired from previous data.
 This composite objective function can be formally expressed as:
-$\mathcal{L}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{MEM}}.$ (2)
-As a canonical example of restriction-based methods, we examine logit-based distillation, which ensures consistency between the logit distributions produced by the old and new models for known classes.
-Specifically, the restriction term $\mathcal{L}_{\text{MEM}}$ is defined as a distillation loss $\mathcal{L}_{\text{KD}}$.
-Let $\hat{q}(\boldsymbol{x})$ and $q(\boldsymbol{x})$ represent the logits of the teacher and student models.
+$\mathcal{L}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{KD}}.$ (2)
+As a canonical example of distillation-based methods, we examine logit-based distillation, which ensures consistency between the logit distributions produced by the old and new models for known classes.
+Specifically, let $\hat{q}(\boldsymbol{x})$ and $q(\boldsymbol{x})$ represent the logits of the teacher and student models.
 The cross-entropy version of $\mathcal{L}_{\text{KD}}$ is defined as:
@@
 where the $\tau$ is the temperature scalar.
-Moreover, most methods are complemented by a replay buffer that stores a subset of representative exemplars from previous tasks.
-These exemplars are interleaved with data from the current task during training to help maintain performance on past tasks.
-Specifically, when learning the $t$-th task, the model is conventionally trained only on the new dataset $D_{t}$.
-With the replay buffer, the training dataset is supplemented with $D_{\text{exp}}$ retrieved from the buffer, forming a combined training set $D_{\text{train}}=D_{t}\cup D_{\text{exp}}$.
-Although the model’s parameters are dynamically updated throughout the incremental learning process, the use of knowledge distillation and the replay buffer ensures that a portion of knowledge from previous tasks remains preserved.
-Replay ensures the model continues to minimize the empirical risk on a representative subset of past tasks, anchoring the parameters in regions consistent with previous data.
+Most of these methods are complemented by a replay buffer that stores a subset of representative exemplars from previous data.
+These exemplars are interleaved with data from the current stage during training to help maintain performance on past data.
+Specifically, in the conventional setting, the model is trained only on the newly arrived data $D_{t}$.
+With the replay buffer, the training data is supplemented with $D_{\text{exp}}$ retrieved from the buffer, forming a combined training set $D_{\text{train}}=D_{t}\cup D_{\text{exp}}$.
+Although the model’s parameters are dynamically updated throughout the continual learning process, the use of knowledge distillation and the replay buffer ensures that a portion of knowledge from previous data remains preserved.
+Replay ensures the model continues to minimize the empirical risk on a representative subset of past data, anchoring the parameters in regions consistent with previous data.
 Concurrently, knowledge distillation functions as a functional regularizer, ensuring that the mappings for old classes remain stable by aligning the output logits (or distributions) with those of the prior model.
 Consequently, a model trained with replay and distillation preserves substantial knowledge of past classes, resulting in a single unified model that is broadly effective but suboptimal.
-We posit that the knowledge stored in the model can be efficiently activated with minor task-specific parameters.
+We posit that the knowledge stored in the model can be efficiently activated with minor auxiliary parameters.
 4 The Proposed Method
@@
 2.
-We treat a replay and distillation-trained network as a base model that accumulates knowledge across tasks but remains suboptimal in accuracy, and then augment it with task-specific plugins that inject residual knowledge into the extracted features to improve performance.
-In this work, each plugin is instantiated with a LoRA-style low-rank parameterization, but is used as a lightweight task extension rather than a conventional fine-tuning update.
+We treat a distillation-based CL network as the base model that accumulates knowledge across each CL stage but remains suboptimal in accuracy, and then augment it with auxiliary plugins that inject residual knowledge into the extracted features to improve performance.
 The proposed DLC is detailed in the following subsections.
-4.1 Task-Specific Plugins Deployment
-Parameter-efficient tuning is typically used in transfer learning by adapting a feature extractor with frozen parameters, where layer inputs remain largely stationary.
-Here, we revisit it beyond this static-backbone assumption and study it under replay- and distillation-based class-incremental learning, where old-task outputs are constrained, and intermediate feature drift is therefore controlled during continual updates.
-Concretely, after training on task $t$, the output deviation at layer $\ell$ is bounded by $K_{\ell}\Gamma_{t}$ (see Appendix), where $K_{\ell}$ depends on hyperparameters such as network structure and distillation temperature, and $\Gamma_{t}$ reflects the distribution discrepancy induced by the chosen replay and distillation strategy.
-This bound directly limits the worst-case feature drift on the old-task distribution, enabling stable optimization of lightweight auxiliary parameters in this dynamic regime.
-We introduce a plugin extension design and maintain an independent plugin set for each task.
-Specifically, for each task $t$, we create a dedicated set of $k$ plugins for $\phi(\boldsymbol{x})$, denoted as $L_{t}$.
-The $\phi(\boldsymbol{x})$ equipped with $L_{t}$ is denoted by $\phi^{L_{t}}(\boldsymbol{x})$.
-In DLC, plugins are deployed in the final layers of $\phi(\boldsymbol{x})$.
-Taking a CNN-based architecture as an example, the first plugin in $L_{t}$ is attached to the last convolutional layer of $\phi(\boldsymbol{x})$, and the remaining plugins are placed on progressively earlier layers moving backward through the network.
-This design is motivated by two considerations.
-First, deeper features are more semantically meaningful and are directly used for classification, so injecting task-specific residuals at this level can more effectively enhance inter-class discriminability.
-Second, it is more compatible with distillation-based objectives.
-If plugins are deployed in shallow layers, residual changes in low-level features would propagate through the network and amplify logit deviations, increasing the optimization difficulty of distillation losses that rely on output alignment between old and new models.
-In DLC, the plugins and $\phi(\boldsymbol{x})$ undergo a decoupled two-phase training procedure.
-Specifically, when a new task $t$ arrives, we first train the base model following the underlying baseline procedure, updating only $\phi(\boldsymbol{x})$ and the classifier $\boldsymbol{W}$ while keeping all plugins frozen.
-After this phase, $\phi(\boldsymbol{x})$ is frozen to prevent knowledge drift, and $L_{t}$ is trained with the same data.
-Upon completion, $L_{t}$ is permanently frozen to safeguard the acquired task-specific knowledge.
-The training objective for all plugins incorporates the cross-entropy loss and an auxiliary loss [33] commonly used in expansion-style methods, formulated as:
+4.1 Classifier-Proximal Plugin Deployment
+We introduce a plugin extension design for continual learning by maintaining an independent residual plugin for each learning stage.
+Specifically, for stage $t$, we attach a dedicated plugin $L_{t}$ for the feature extractor $\phi(\boldsymbol{x})$, and denote the resulting plugin-enhanced representation by $\phi^{L_{t}}(\boldsymbol{x})$.
+Rather than modifying the backbone weights directly, each plugin acts as a lightweight residual branch that performs stage-specific semantic correction on the extracted features.
+In this way, DLC complements the continually updated backbone with additional discriminative knowledge in a plug-and-play manner, while preserving the backbone as the main carrier of shared knowledge across stages.
+DLC adopts a classifier-proximal deployment strategy, where plugins are attached to the final layer of the $\phi(\boldsymbol{x})$, i.e., the layer closest to the classifier.
+This placement is beneficial for two reasons.
+First, classifier-proximal features are already semantically mature and directly support classification, so residual correction at this stage can more effectively improve inter-class discriminability.
+Second, the injected perturbation is localized near the classifier and thus minimally disrupts the overall feature extraction process.
+Compared with shallow-layer intervention, such a design is more compatible with distillation-based objectives, since shallow-layer feature changes would otherwise propagate through the network and amplify output deviations, making output alignment between the old and current models harder to optimize.
+The plugins and $\phi(\boldsymbol{x})$ are trained in a decoupled two-phase manner.
+When a new task $D_{t}$ arrives, we first follow the underlying baseline to train the backbone and classifier while keeping all plugins frozen.
+After this phase, the backbone is frozen, and only the newly introduced plugin $L_{t}$ is optimized on the same data.
+Once trained, $L_{t}$ is permanently frozen to preserve the acquired stage-specific knowledge.
+The plugin objective combines the cross-entropy loss with an auxiliary loss commonly adopted in expansion-style methods [38]:
 $\mathcal{L}_{\text{plg}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}.$ (5)
-The two-phase procedure ensures that plugins are trained without interference from distillation losses, enabling specialized learning for the target task.
-Since $\phi(\boldsymbol{x})$ remains frozen during plugin updates, plugin training is stable, and it also prevents the baseline training of $\phi(\boldsymbol{x})$ from being affected by plugin training, preserving the plug-and-play property of DLC.
-During inference, all plugins are activated to extract their respective feature representations.
-These representations are then concatenated to form an integrated feature vector, which enhances the base model through a multi-task feature ensemble.
-For a linear classifier $\boldsymbol{W}$, the classification process can be formulated as:
-$\hat{y}=\boldsymbol{W}[\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),...,\phi^{L_{T}}(\boldsymbol{x})]$ (6)
-where $T$ denotes the maximum index of known tasks.
-Notably, we do not need to run the full extractor $T$ times: each plugin injects a residual only at its deployment layer, so the backbone activations up to the earliest deployed layer are shared and computed once, avoiding redundant computation and improving inference efficiency.
-4.2 LoRA-Style Residual Plugin
-To instantiate each task-specific plugin with minimal parameters, we adopt a LoRA [13]-style low-rank residual parameterization.
-This design is well-suited for plugin deployment: the rank $r$ provides a simple capacity knob under a fixed budget, the same low-rank residual form applies to both linear and convolutional operators, and the residual structure supports our decoupled training and inference aggregation since plugins can be independently trained, frozen, and combined without modifying the backbone.
-We emphasize that the backbone weights are kept intact throughout, and the low-rank component is introduced as an auxiliary module that adds a task-specific residual to the feature.
-That is, the LoRA plugin does not replace or factorize the original weights, but injects additional task knowledge via an additive low-rank branch, which aligns with our plug-and-play deployment setting.
-Since most CIL pipelines train CNN-based backbones from scratch, we instantiate the plugin for convolutional operators using a LoRA-style low-rank residual branch (implemented as two convolutions).
+This decoupled procedure allows plugins to learn stage-specific corrections without being directly entangled with distillation optimization, while also preventing plugin training from interfering with the baseline update of the backbone.
+During inference, all learned plugins are activated to produce their corresponding plugin-enhanced representations.
+Because each plugin encodes a different stage-specific residual refinement on top of the shared backbone output, these representations provide diverse semantic cues with potentially complementary information.
+Their aggregation can thus be viewed as a classifier-proximal form of multi-stage feature reuse, in spirit similar to DenseNet [15] feature reuse, enriching the final prediction with knowledge accumulated across learning stages.
+For a linear classifier $\boldsymbol{W}$, the classification process can be written as:
+$\hat{y}=\boldsymbol{W}[\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),\ldots,\phi^{L_{T}}(\boldsymbol{x})],$ (6)
+where $T$ denotes the total number of learned stages so far.
+Notably, the full feature extractor does not need to be executed $T$ times: since each plugin injects a residual after the last layer, the backbone activations before the deployed layer can be shared and computed once, avoiding redundant computation.
+For continual learning to remain practical in real-world scenarios, the goal is not to allocate an extractor-level or layer-level plugin for every learning stage, but to introduce a lightweight auxiliary module plugin with minimal storage overhead.
+Such a design naturally calls for parameter-efficient fine-tuning (PEFT) modules [11].
+However, these modules have previously been studied mostly in transfer learning, where they are typically deployed on top of a knowledge-rich pretrained backbone.
+Distillation-based continual learning presents a fundamentally different regime: the backbone is not a static pretrained model with abundant prior knowledge, but a continually updated learner whose knowledge is progressively accumulated and preserved through replay and distillation.
+As discussed in Sec.
+3.2, knowledge distillation together with a replay buffer helps retain previously acquired knowledge, thereby constraining distribution drift during continual updates.
+This makes it possible to revisit PEFT modules in a dynamic backbone.
+Concretely, after training on stage $t$, the output deviation at layer $\ell$ is bounded by $K_{\ell}\Gamma_{t}$, where $K_{\ell}$ depends on hyperparameters such as the network architecture and distillation temperature, and $\Gamma_{t}$ reflects the distribution discrepancy induced by the adopted replay and distillation strategy.
+This bound directly limits the worst-case feature drift on the old-stage distribution, enabling stable optimization of PEFT modules in this dynamic regime.
+A detailed proof is provided in the supplementary material.
+4.2 LoRA-Style Plugins
+To instantiate each stage-specific plugin with minimal parameters, we adopt a LoRA [14]-style low-rank residual parameterization.
+This design is well-suited for plugin deployment: the residual structure supports our decoupled training and inference aggregation since plugins can be independently trained, frozen, and combined without modifying the backbone.
+We emphasize that the backbone weights are kept intact throughout, and the low-rank component is introduced as an auxiliary module that adds residual to the feature.
+That is, the LoRA plugin does not replace or factorize the original weights, but injects additional knowledge via an additive low-rank branch, which aligns with our plug-and-play deployment setting.
+Following the classifier-proximal deployment strategy in Sec.
+4.1, this plugin is attached to the final feature layer of $\phi(\boldsymbol{x})$; for CNN backbones, it is instantiated on the representation produced by the last convolutional layer.
+However, LoRA was originally proposed for Transformer models, while distillation-based CL methods are often built upon CNN-based backbones trained from scratch.
+We therefore provide a convolutional LoRA-style plugin instantiation for this setting, implemented as a lightweight low-rank residual branch with two convolutions.
 For a convolutional layer with weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times K\times K}$ and input $\boldsymbol{x}\in\mathbb{R}^{N\times C_{\text{in}}\times H\times W}$, the original forward pass is $\boldsymbol{h}=W_{\text{conv}}\ast\boldsymbol{x}$, where $\ast$ denotes the convolution operation.
@@
 $\boldsymbol{h}^{\prime}=W_{\text{conv}}\ast\boldsymbol{x}+\frac{\alpha}{r}(\mathbf{B}\ast(\mathbf{A}\ast\boldsymbol{x})),$ (9)
-where $\alpha$ is a scaling hyperparameter, and $r$ is the rank.
-Compared to the model expansion, instantiating task-specific plugins with a LoRA-style parameterization can bring significant parameter advantages.
+where $\alpha$ is a scaling hyperparameter, and $r$ is the rank. Following [12], $r$ and $\alpha$ are set to the recommended values of 8 and 16.
+Compared to the model or layer-level expansion, instantiating the LoRA-style PEFT plugins can bring significant parameter advantages.
 Specifically, for a convolutional layer ($C_{in}$, $C_{out}$, $K\times K$), the space complexity of its parameters is $\Theta(C_{in}C_{out}K^{2})$, while the plugin only occupies $\Theta(r(C_{in}K^{2}+C_{out}))$.
-For expanding one layer of CNN over $T$ tasks, considering that in a CNN in CIL, $C_{out}$ is generally greater than or equal to $C_{in}$, we set $C_{in}=C_{out}=C$, and $K$ can be regarded as a constant (usually 3 or 7).
-Since $r$ is low-rank, it can also be regarded as a constant (generally set to 8 or 16).
+For expanding one layer of CNN over $T$ stages, considering that in a CNN in CL, $C_{out}$ is generally greater than or equal to $C_{in}$, we set $C_{in}=C_{out}=C$, and $K$ can be regarded as a constant (usually 3 or 7).
+Moreover, since the low rank $r$ is set to 8, it can be treated as a constant.
 Then, using LoRA-style plugin instantiation reduces the parameter-space complexity of expanding a single CNN layer from $\Theta(TC^{2})$ to $\Theta(TC)$.
@@
 4.3 Dynamic Weighting Unit
-For a sample from task $t$, all the plugins trained on other tasks are considered non-target plugins.
-Given that the replay buffer contains only a minimal fraction of the original data for each task, these non-target plugins are not trained on the full data from task $t$.
-In particular, plugins trained on tasks 1 to t1t-1 never encounter samples from the task $t$.
-Consequently, the residuals extracted by these plugins could be largely irrelevant to the task $t$, thus introducing uninformative features or even detrimental noise to the classifier.
-To mitigate the influence of uninformative residuals from non-target plugins, we will create a weighting unit corresponding to concentrated representations when training on a new task.
-This unit comprises two linear layers followed by a sigmoid function $\sigma(\cdot)$, denoted as $\omega(\mathbf{x})=\sigma\Big(\mathbf{W}_{2}\text{ReLU}\big(\mathbf{W}_{1}\mathbf{x}\big)\Big),\mathbf{x}\in\mathbb{R}^{k},\mathbf{W}_{1}\in\mathbb{R}^{k\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times k}$, where the sigmoid function normalizes the output to the [0, 1] range, indicating their importance weights.
-The concatenated representations are passed through this unit before being fed into the classifier, then weighted representations is derived from $\tilde{\mathbf{h}}^{(i)}=\mathbf{\omega}^{(i)}\odot\mathbf{h}^{(i)}$, where $\odot$ is element-wise multiplication operator.
+For any sample from stage $t$, all the plugins trained on other tasks $D_{i\neq t}$ are considered non-target plugins.
+Given that the replay buffer contains only a minimal fraction of the original data, these non-target plugins are not trained on the full $D_{t}$.
+In particular, plugins trained on stages 1 to t1t-1 never encounter samples from $D_{t}$.
+Consequently, the residuals extracted by these plugins could be largely irrelevant to the stage $t$, thus introducing uninformative features or even detrimental noise to the classifier.
+To mitigate their influence, we will create a weighting unit corresponding to concentrated representations when training on a new stage.
+This unit comprises two linear layers followed by a sigmoid function $\sigma(\cdot)$, denoted as $\omega(\mathbf{x})=\sigma\Big(\mathbf{W}_{2}\text{ReLU}\big(\mathbf{W}_{1}\mathbf{x}\big)\Big),\mathbf{x}\in\mathbb{R}^{k},\mathbf{W}_{1}\in\mathbb{R}^{k\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times k}$, where $\sigma(\cdot)$ normalizes the output to the [0, 1] range.
+During inference, we first compute the importance weights for the aggregated plugin-enhanced representation:
+$\boldsymbol{\omega}=\omega\!\left([\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),\ldots,\phi^{L_{T}}(\boldsymbol{x})]\right),$ (10)
+then perform inference using the reweighted representation:
+$\hat{y}=\boldsymbol{W}\left(\boldsymbol{\omega}\odot[\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),\ldots,\phi^{L_{T}}(\boldsymbol{x})]\right).$ (11)
+Where the $\odot$ denotes element-wise multiplication.
 The weighting unit can be regarded as a gate for the classifier.
 Jointly trained with the classifier, it helps the model adapt to the high-dimensional feature space and automatically assigns lower weights to uninformative or noisy features.
-The automatically trained weighting unit fails to explicitly model the varying importance of representations from different tasks.
-For a sample $i$ in task $t$, we denote the representations adapted by LoRA plugins trained on tasks 1 to t1t-1 as $\mathbf{h}^{(i)}_{\text{pre}}$ and those adapted by the remaining LoRA plugins as $\mathbf{h}^{(i)}_{\text{pos}}$, respectively.
-Then $\mathbf{h}^{(i)}_{\text{pre}}$ should be assigned lower weights as corresponding LoRA plugins were not optimized on task $t$.
-However, since the ground-truth task index of each sample is unavailable during the CIL testing, we cannot explicitly control the weighting unit based on task identity.
+The automatically trained weighting unit fails to explicitly model the varying importance of representations from different stages.
+For a sample $i$ in stage $t$, we denote the representations adapted by plugins trained on stages 1 to t1t-1 ($\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),\ldots,\phi^{L_{t-1}}(\boldsymbol{x})]$) as $\mathbf{h}^{(i)}_{\text{pre}}$ and those adapted by the remaining plugins as $\mathbf{h}^{(i)}_{\text{pos}}$, respectively.
+Then $\mathbf{h}^{(i)}_{\text{pre}}$ should be assigned lower weights as corresponding plugins were not optimized with samples in $D_{t}$.
+Since the stage of each sample is unavailable during the CL testing, we cannot explicitly control the weighting unit based on stage identity.
 Instead, we employ a supervised regression training to guide the weighting unit to approximate the desired weighting scheme.
-We decompose the weight $\mathbf{\omega}^{(i)}$ for sample $i$ into $\mathbf{\omega}^{(i)}_{\text{pre}},\mathbf{\omega}^{(i)}_{\text{pos}}$] Then the decision process of the linear classifier can be rewritten as follows:
-$\boldsymbol{q}^{i}=\mathbf{W}_{\text{pre}}(\mathbf{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})+\mathbf{W}_{\text{pos}}(\mathbf{\omega}^{(i)}_{\text{pos}}\odot\mathbf{h}^{(i)}_{\text{pos}})$.
+We decompose the weight $\boldsymbol{\omega}^{(i)}$ for sample $i$ into $\boldsymbol{\omega}^{(i)}_{\text{pre}},\boldsymbol{\omega}^{(i)}_{\text{pos}}$] Then the decision process of the linear classifier can be rewritten as follows:
+$\boldsymbol{q}^{i}=\mathbf{W}_{\text{pre}}(\boldsymbol{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})+\mathbf{W}_{\text{pos}}(\boldsymbol{\omega}^{(i)}_{\text{pos}}\odot\mathbf{h}^{(i)}_{\text{pos}})$.
 (12)
-Formally, the contribution of $\mathbf{h}^{(i)}_{\text{pre}}$ to the classifier’s logits can be expressed as $\mathbf{\xi}=\mathbf{W}_{\text{pre}}(\mathbf{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})$.
+Formally, the contribution of $\mathbf{h}^{(i)}_{\text{pre}}$ to the classifier’s logits can be expressed as $\mathbf{\xi}=\mathbf{W}_{\text{pre}}(\boldsymbol{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})$.
 To generate effective weights, an importance-aware optimization objective could be incorporated into the training process, denoted as $\mathcal{L}_{\text{IA}}$.
 This loss function employs the Mean Squared Error (MSE), expressed as follows:
-$\mathcal{L}_{\text{IA}}=\frac{1}{k}{\|\mathbf{\omega}-\mathbf{\omega}_{\text{ideal}}\|}_{2},$ (11)
-where the $\mathbf{\omega}_{\text{ideal}}$ is the ideal weight output, it has $\mathbf{\omega}^{(i)}_{\text{ideal-pos}}=1$ and $\mathbf{\omega}^{(i)}_{\text{ideal-pre}}=0$.
-In the ideal training case, the model has $\mathbf{\omega}^{(i)}_{\text{pos}}\to 1$ and $\mathbf{\omega}^{(i)}_{\text{pre}}\to 0$, then it has $\mathbb{E}[\mathbf{\xi}]\to 0$ and $\text{Var}[\mathbf{\xi}]\to 0$, so the $\mathbf{h}^{(i)}_{\text{pre}}$ have negligible influence on prediction logits.
+$\mathcal{L}_{\text{IA}}=\frac{1}{k}{\|\boldsymbol{\omega}-\boldsymbol{\omega}_{\text{ideal}}\|}_{2},$ (13)
+where the $\boldsymbol{\omega}_{\text{ideal}}$ is the ideal weight output, it has $\boldsymbol{\omega}^{(i)}_{\text{ideal-pos}}=1$ and $\boldsymbol{\omega}^{(i)}_{\text{ideal-pre}}=0$.
+In the ideal training case, the model has $\boldsymbol{\omega}^{(i)}_{\text{pos}}\to 1$ and $\boldsymbol{\omega}^{(i)}_{\text{pre}}\to 0$, then it has $\mathbb{E}[\mathbf{\xi}]\to 0$ and $\text{Var}[\mathbf{\xi}]\to 0$, so the $\mathbf{h}^{(i)}_{\text{pre}}$ have negligible influence on prediction logits.
 Even if the ideal scenario is hard to achieve, the classifier can still produce more accurate predictions as long as the representations adapted by reliable plugins are assigned higher weights, while those adapted by non-target plugins are appropriately suppressed.
@@
 Table 1: Last performance $A_{T}$ and average performance $\overline{A}$ comparison.
-‘-DLC’ indicates the method using the DLC framework.
-‘Extra/Feat #P’ denotes the total number of extra parameters required for deploying DLC and the feature extractor, respectively.
 Methods CIFAR-100 Tiny-ImageNet ImageNet-100
@@
 Extra/Feat #P (MB) 0.15 0.46 0.07 0.46 0.04 0.46 0.45 11.17 0.52 11.17
+Figure 3: Comparison of $A_{T}$ with/without DLC when enhanced with AutoAugment (AA) , $\mathcal{L}_{\text{inter-intra}}$ and C-flat.
 5 Experiments
@@
 Sec.
-5.3 investigates the performance of SOTA CIL methods under the memory-aligned case.
+5.3 further evaluates DLC under challenging CL settings, including long-sequence, few-shot, and class-imbalanced scenarios.
 Sec.
-5.4 evaluates the compatibility of DLC with other data augmentation and plug-and-play loss functions for CIL.
+5.4 investigates the performance of SOTA CL methods under the memory-aligned case.
 Sec.
@@
 Datasets and Split.
-We chose three benchmark datasets commonly used in CIL, including CIFAR-100 [15], ImageNet-100 [9], and Tiny-ImageNet-200 [28], where the number following the dataset is the total number of classes included.
-Following the ‘B/Base-m, Inc-n’ rule proposed by [7], we slice each of the above three datasets into CIFAR-100 B10 Inc10, CIFAR-100 B50 Inc10, CIFAR-100 B5 Inc5, ImageNet-100 B50 Inc10, and Tiny-ImageNet B40 Inc40 for the experiment, where $m$ is the number of classes in the first incremental task, and $n$ represents that of every remaining stage.
+We chose three benchmark datasets commonly used in CL, including CIFAR-100 [17], Tiny-ImageNet-200 [32], and the large-scale ImageNet-100 [7].
+Following the ‘B/Base-m, Inc-n’ rule proposed by [5], we slice each of the above three datasets into CIFAR-100 B10 Inc10, CIFAR-100 B50 Inc10, CIFAR-100 B5 Inc5, ImageNet-100 B50 Inc10, and Tiny-ImageNet B40 Inc40 for the experiment, where $m$ is the number of classes in the first learning stage, and $n$ represents that of every remaining stage.
 For a fair comparison, we ensure that the training and test datasets are identical for all methods.
-Detailed information about the dataset is placed in the Appendix section.
+Detailed information about the dataset is provided in the Appendix.
 Implementation Details.
@@
 We employ the lightweight ResNet-32 tailored for CIFAR datasets on CIFAR-100, while utilizing ResNet-18 for all other benchmark datasets.
-The optimizer employs SGD with a fixed learning rate of 0.1, incorporating a multi-step schedule.
+The optimizer employs SGD incorporating the multi-step schedule, with a learning rate of 0.1.
 The detailed training hyperparameters, like the optimizer and batch size, largely follow the default settings in PyCIL.
-All the methods are run using the same random number seed.
+All the methods are run using the same random seed.
 To eliminate potential confounding effects from increased training epochs, all baselines are run under both the PyCIL-recommended epoch setting and the total epoch budget of DLC’s two-stage training, and the maximum accuracy is reported.
 Evaluation Metrics.
-Following the benchmark protocol settings [23], we use $A_{b}$ to denote the $b$-stage accuracy on the test set that includes all known classes after trained with $D_{1},D_{2},\cdots,D_{b}$, $\overline{A}=\frac{1}{T}\sum_{b=1}^{T}A_{b}$ is average accuracy over $T$ tasks, and $A_{T}$ is the last accuracy on the test set that includes all learned tasks.
+Following the benchmark protocol settings [27], we use $A_{b}$ to denote the $b$-stage accuracy on the test set that includes all known classes after trained with $D_{1},D_{2},\cdots,D_{b}$, $\overline{A}=\frac{1}{T}\sum_{b=1}^{T}A_{b}$ is average accuracy over $T$ stages, and $A_{T}$ is the last accuracy on the test set that includes all learned stages.
 We use $\overline{A}$ and $A_{T}$ to evaluate the model’s performance, which reflects the model’s dynamic learning capacity and generalization ability, respectively.
-5.2 Effectiveness on Distillation-Based CIL
-In this section, we evaluate the performance improvement brought by the proposed DLC framework across different distillation-based CIL objectives, including iCaRL [23], WA [38], BiC [31], DAKD [12], and MKD [20].
+5.2 Effectiveness on Distillation-Based CL
+In this section, we evaluate the performance improvement brought by the proposed DLC framework across different distillation-based CL objectives, including iCaRL [27], WA [43], BiC [36], DAKD [13], and MKD [24].
 Additionally, a simple herding-based Replay method [27] is included to examine DLC’s contribution in scenarios relying solely on the rehearsal without distillation.
-We report the $\overline{A}$ and $A_{T}$ of these established methods with and without DLC across three datasets under varying task configurations.
+We report the $\overline{A}$ and $A_{T}$ of these established methods with and without DLC across three datasets under varying configurations.
 In all experiments, the replay buffer size is fixed at 2,000 exemplars.
-DLC expands only one LoRA plugin per task, applied to the last convolutional layer of the feature extractor.
+DLC expands only one LoRA plugin per stage, i.e., each plugin consists of a single pair of low-rank matrices $\mathbf{A}$ and $\mathbf{B}$.
 We also report the total parameters required by DLC and the size of the entire feature extractor for comparison.
 The results are summarized in Table 1.
-The results show that DLC significantly improves all classical replay or distillation-based methods with only one LoRA plugin per task.
+The results show that DLC significantly improves all distillation-based methods.
 Moreover, the parameter overhead introduced by DLC is negligible.
-DLC achieves up to 8% accuracy improvement while expanding only 4.8% to 32% parameters equivalent to the feature extractor, highlighting its practical value in CIL.
+DLC achieves up to 8% accuracy improvement while expanding only 4.8% to 32% parameters equivalent to the feature extractor, highlighting its practical value in CL.
 Fig.
 4 shows the confusion matrix heatmap of iCaRL and iCaRL-DLC as an example.
-DLC increases the thermal intensity along the diagonal for most classes, demonstrating that it could enhance the model’s accuracy across most tasks.
-In addition, we compare the final accuracy of DLC-enhanced iCaRL, WA, and BiC with and without the weighting unit, as shown in Fig.
-4.
-Results indicate that the proposed weighting unit contributes a certain accuracy gain.
+DLC increases the thermal intensity along the diagonal for most classes, demonstrating that it could enhance the model’s accuracy across most classes.
 Figure 4: Confusion matrix heatmaps of iCaRL with and w/o DLC on CIFAR-100 B10 Inc10.
 Figure 5: Comparison of $A_{T}$ with and without the weighting unit for DLC-enhanced methods on CIFAR-100.
+Figure 6: Comparison of $A_{T}$ with plugins deployed at different positions (FL/ML/CPL) on CIFAR-100.
+In addition to DLC, other plug-and-play enhancement strategies have been proposed to enhance CL methods.
+For example, some works employ the data augmentation strategy AutoAugment for continual learning, particularly on the CIFAR-100 dataset [33, 44].
+Several plug-and-play loss functions have also been introduced, including $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ for distillation methods [8], as well as C-flat [18].
+We evaluate the compatibility of DLC with such enhanced strategies.
+Specifically, we compare the $A_{T}$ of baseline methods using AutoAugment, the loss $\mathcal{L}_{\text{inter-intra}}$, and C-flat with and without DLC.
+For experiments with $\mathcal{L}_{\text{inter-intra}}$, we follow the original implementation and use ResNet-18 as the backbone.
+As shown in the Fig.
+3, DLC consistently brings significant improvements, demonstrating its compatibility with other plug-and-play enhancements.
+Moreover, we conduct two ablation studies.
+First, we evaluate the effect of the proposed weighting unit by comparing the final accuracy of DLC-enhanced iCaRL, WA, and BiC with and without it, as shown in Fig.
+5.
+Results indicate that the weighting unit contributes to improved performance.
+Second, we investigate the importance of classifier-proximal deployment by placing the plugin at different depths of the feature extractor, including the first layer (FL), middle layer (ML), and the classifier-proximal layer (CPL).
+As shown in Fig.
+6, only classifier-proximal deployment improves performance, whereas earlier deployment causes increasingly severe degradation, indicating poorer distillation compatibility.
+Figure 7: Comparison of $A_{T}$ with/without DLC under challenging continual learning settings, including long-sequence, few-shot, and class-imbalanced scenarios.
+5.3 Evaluation under Challenging CL Settings
+Distillation-based CL methods are sometimes deployed in extreme scenarios, such as long learning sequences, few-shot, and class-imbalanced data.
+We therefore further evaluate DLC under these settings.
+For the long-sequence setting, we use Tiny-ImageNet with the B5Inc5 protocol, resulting in 40 stages.
+For the few-shot setting, we follow the protocol in [16] on CIFAR with B60Inc5, where each class contains only 5 samples after the second stage.
+For the class-imbalance setting, following Liu et al.
+[21], we adopt the CIFAR-100 B50Inc5 protocol with an imbalance ratio $\rho=0.01$ (i.e., head-to-tail ratio takes 100).
+The results are shown in Fig.
+7.
+DLC consistently improves distillation-based CL methods even under these challenging settings.
+Figure 8: Comparison of inference FLOPs and per-sample inference latency with and w/o DLC.
 5.4 Comparison under Fixed Memory Budget
 Table 2: Performance comparison on CIFAR-100 with aligned memory cost.
-‘#P’ represents the number of parameters (million).
-$\varepsilon$’ denotes the number of exemplars, and ‘MS’ denotes the memory size (MB).
-All methods use ResNet-32 as the backbone.
+‘#P’ represents the number of model parameters (million).
+$\varepsilon$’ denotes the number of exemplars, and ‘MS’ denotes the total memory size (MB).
 Method CIFAR100 B10 Inc10
@@
 WA-DLC 23.5 0.54 7327 75.02 64.79
-Prior research [42] has demonstrated that model-level expansion-based methods achieve state-of-the-art performance across all standard CIL benchmarks.
-However, these approaches overlook the substantial memory overhead induced by dynamic architecture growth.
-Such methods implicitly introduce additional memory budgets, i.e., a model buffer for retaining parameters of the expanded model, which grants them a considerable and unaccounted advantage over other methods that do not store expansion models or adopt more lightweight expansion strategies.
-Following the protocol established in [43], this section presents a memory-aligned comparison of CIL baselines.
+Prior studies have shown that model-level expansion methods can achieve state-of-the-art performance in continual learning, but their gains often rely on substantial and uncounted memory overhead for storing expanded model parameters, making direct comparison with distillation-based methods unfair.
+To further demonstrate the parameter efficiency and accuracy of DLC under a fair budget constraint, we follow the protocol established in [48] and present a memory-aligned comparison of CL baselines.
 Under a fixed 23.5 MB memory budget, the memory not consumed by a method’s parameters is allocated to replay exemplars.
@@
 Among them, DER, MEMO, and TagFex have incorporated the weight alignment from WA in their implementations.
-As summarized in Table 2, deploying DLC to the distillation-based method WA under the memory-aligned setting yields the highest classification accuracy.
-5.4 Compatibility with other Plug-and-Play Enhancements
-Figure 5: Comparison of $A_{T}$ with/without DLC on CIFAR-100 when enhanced with AutoAugment (AA) and $\mathcal{L}_{\text{inter-intra}}$.
-In addition to DLC, other plug-and-play strategies have been proposed to enhance CIL methods.
-For example, the data augmentation technique AutoAugment [6] has been employed in the CIFAR-100 dataset by [29, 39].
-Meanwhile, two additional training losses $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ specifically designed for logit-based distillation methods have been introduced in [10].
-This subsection evaluates the compatibility of DLC with such enhanced strategies.
-Specifically, we compare the accuracy of methods using AutoAugment and the loss $\mathcal{L}_{\text{inter-intra}}$ with and without DLC.
-For experiments with $\mathcal{L}_{\text{inter-intra}}$, we follow the original implementation and use ResNet-18 as the backbone.
-As shown in the Fig.
-5, DLC consistently brings significant improvements, demonstrating its compatibility with other plug-and-play enhancements.
-5.5 Additional Computational Overhead from DLC
+All methods use the same ResNet backbone.
+As summarized in Table 2, WA equipped with DLC achieves the highest classification accuracy under the memory-aligned setting.
+This highlights that DLC delivers a superior memory-accuracy trade-off.
+5.5 Additional Computational Overhead
 This section investigates the additional inference cost introduced by deploying DLC.
-Using iCaRL as a representative baseline, we evaluate two deployment variants: iCaRL equipped with one LoRA plugin per task, and iCaRL equipped with both the LoRA plugins and the weighting unit.
+Using iCaRL as a representative baseline, we evaluate two deployment variants: iCaRL equipped with only DLC plugins, and iCaRL equipped with both the plugins and the weighting unit.
 Concretely, on ImageNet-100, we report the total FLOPs and the average inference latency per sample (ms) measured on the same test set.
 The results are shown in Fig.
-6.
-Overall, neither FLOPs nor latency increases substantially as the number of tasks grows.
-Instead, both grow mildly as the task sequence extends.
+8.
+Overall, neither FLOPs nor latency increases substantially as the number of stages grows.
+Instead, both grow mildly as the stage sequence extends.
 Even under long sequences such as B5 Inc5, the increases in FLOPs and inference time are only about 4%.
-These results indicate that DLC provides a lightweight and computation-friendly parameter extension for enhancing distillation-based CIL.
-Figure 6: Comparison of inference FLOPs and per-sample inference latency with and w/o DLC under different task settings.
+These results indicate that DLC provides a lightweight and computation-friendly parameter extension for enhancing distillation-based CL.
 6 Conclusion
-In this work, we address a challenge in distillation-based class-incremental learning: improving the model performance with only a small amount of task-specific auxiliary parameters, without breaking the single-model distillation setting.
-We propose DLC, a plug-and-play plugin extension framework that augments replay and distillation-based baselines with task-specific LoRA-style residual plugins.
-DLC trains each plugin in a decoupled manner and then freezes it, which both avoids interference from distillation objectives and preserves the acquired task knowledge.
-We further introduce a lightweight weighting unit to down-weight irrelevant plugin features at inference.
-Extensive experiments demonstrate that DLC consistently improves replay and distillation-based CIL baselines with few parameter overhead.
+In this work, we address a challenge in distillation-based continual learning: improving the model performance with only a small amount of auxiliary parameters, without breaking the single-model distillation setting.
+We propose DLC, a plug-and-play plugin extension framework that augments replay and distillation-based baselines with LoRA-style residual plugins.
+DLC deploys each plugin at the classifier-proximal layer and trains it in a decoupled manner, allowing residual correction to directly refine features while remaining compatible with distillation.
+We further introduce a lightweight weighting unit to down-weight irrelevant features in inference.
+Extensive experiments demonstrate that DLC consistently improves distillation-based CL baselines with few parameter overhead.
 Future work will explore more expressive weighting designs for stronger feature integration, as well as earlier-layer plugin deployment to further improve adaptability.
-Appendix
-Appendix A Dataset Description
-CIFAR-100 CIFAR-100 (Canadian Institute for Advanced Research 100-class dataset) is a widely used standard benchmark dataset in the field of computer vision, particularly for image classification tasks.
-It is an expanded version of the CIFAR-10 dataset, curated by the Canadian Institute for Advanced Research.
-A defining feature of this dataset is its fine-grained categorical structure.
-It contains a total of 100 distinct fine-grained classes (fine labels), organized hierarchically under 20 broader superclasses (coarse labels), with each superclass encompassing precisely 5 fine-grained classes.
-This hierarchical organization provides a valuable foundation for researching multi-level classification.
-The dataset comprises 60,000 tiny color images, each with a fixed resolution of 32x32 pixels.
-This compact size makes the dataset manageable for rapid experimentation and algorithm iteration, while simultaneously presenting significant challenges for model recognition capabilities due to the extreme low resolution.
-The 60,000 images are explicitly divided into 50,000 training images and 10,000 test images, meaning each fine-grained class includes 500 training images and 100 test images.
-CIFAR-100 offers broad and diverse category coverage, encompassing common objects (various vehicles, furniture, appliances), animals (mammals, fish, insects), plants (flowers, trees), and natural scenes (mountains, oceans), among others.
-Its extensive number of classes combined with its moderate scale make it an ideal testbed for evaluating and comparing the performance and generalization capabilities of image classification models—especially Convolutional Neural Networks (CNNs)—as well as for exploring advanced tasks like few-shot learning and transfer learning.
-ImageNet-100 ImageNet-100 is a commonly used subset of the well-known large-scale image dataset, ImageNet.
-Its purpose is to significantly reduce computational resource requirements and training time while preserving much of ImageNet’s richness and challenge, making it more suitable for rapid prototyping, algorithm exploration, and research in resource-constrained environments.
-As the name implies, ImageNet-100 consists of 100 classes (synsets) selected from the original ImageNet dataset.
-The selection of these classes is typically not entirely random; it often follows specific strategies (e.g., choosing semantically distinct classes or classes with relatively higher sample counts) to ensure the subset remains representative and diverse.
-The dataset contains approximately 100,000 high-resolution color images curated from these 100 classes.
-This represents a substantial reduction in scale compared to the millions of images in the full ImageNet.
-The standard partition of the dataset usually includes a training set and a validation set.
-The training set contains roughly 1,000 images per class, totaling approximately 100,000 images.
-The validation set contains about 50 images per class, totaling around 5,000 images.
-The images themselves are typically of variable sizes (significantly larger than CIFAR’s 32x32 pixels) and require preprocessing (like cropping and resizing) before being fed into models.
-ImageNet-100 inherits core characteristics from its parent dataset: rich and realistic image content, high semantic distinction between classes, significant intra-class variation (the same object under different viewpoints, lighting, occlusion), and inter-class similarities.
-It is widely used for training and evaluating image classification models, as well as for transfer learning tasks.
-Compared to smaller datasets like CIFAR, models pre-trained on ImageNet-100 generally learn more generalizable visual features suitable for fine-tuning on downstream tasks.
-Compared to the full ImageNet, it offers faster iteration cycles and lower hardware barriers to entry.
-TinyImageNet-200 TinyImageNet-200 is a downsized variant of the ImageNet dataset, specifically created for educational and experimental purposes in deep learning and computer vision research.
-It features a curated subset of 200 distinct object categories (synsets) carefully selected from the full ImageNet hierarchy.
-The dataset contains roughly 100,000 high-resolution color images – significantly fewer than its parent dataset but substantially more than smaller benchmarks like CIFAR.
-Each image in Tiny ImageNet has been consistently downscaled and cropped to a fixed resolution of 64x64 pixels, maintaining a usable size while greatly reducing computational demands compared to processing full-sized ImageNet images.
-The dataset comes with a predefined partition: a training set containing 500 images per class, a separate validation set with 50 images per class, and a test set also containing 50 images per class, totaling approximately 100,000 training images, 10,000 validation images, and 10,000 test images.
-Alongside the images, TinyImageNet provides essential metadata files including a wnids.txt file listing the WordNet IDs for all 200 classes, a words.txt file mapping these IDs to human-readable labels, and a structured val_annotations.txt file listing the ground-truth classes for validation images.
-Its manageable scale, balanced class representation, and standardized structure make Tiny ImageNet-200 a popular choice for academic courses exploring deep learning architectures (particularly CNNs and Vision Transformers), as well as for research prototyping on tasks such as image classification, transfer learning, and model robustness evaluation, especially where computational resources are constrained but larger challenge than CIFAR is needed.
-Appendix B Feature Drift Bounds under Logit Distillation and Herding Replay
-Before applying methods like LoRA to continually updated models, it is imperative to guarantee that the feature extractor’s output deviation for previous tasks remains bounded before and after training.
-Otherwise, if catastrophic forgetting causes a significant shift in outputs, the plugin modules would be unable to extract meaningful residuals.
-Setting.
-Consider a sequence of tasks $D_{1},D_{2},\dots,D_{t}$. At task $t$, the model is trained on
-$D_{\mathrm{train}}=D_{t}\cup D_{\exp},$ (13)
-where $D_{\exp}$ is the class-balanced exemplar set from previous tasks based on the herding replay strategy in iCaRL.
-Let $\theta$ denote the parameters of a convolutional feature extractor $\phi(\boldsymbol{x})$, and let $h_{\ell}(x;\theta)\in\mathbb{R}^{d_{\ell}}$ denote the output of layer $\ell$ in $\phi(\boldsymbol{x})$.
-We denote the parameters of $\phi(\boldsymbol{x})$ as $\theta_{t-1}$ and $\theta_{t}$ for before and after training on task $t$.
-We adopt the logit-based distillation approach as a representative case for analysis.
-To constrain forgetting, the logits are distilled from the teacher to the student.
-The logits of teacher and student are denoted as $\hat{z}(x)$ and $z(x)$.
-For temperature $T>0$, define the softened probabilities
-$\hat{q}^{T}(x)=\mathrm{softmax}\!\big(\hat{z}(x)/T\big),$ (14)
-$q^{T}(x)=\mathrm{softmax}\!\big(z(x)/T\big).$ (15)
-The cross-entropy-based distillation loss is
-$\mathcal{L}_{\mathrm{KD}}(x)=-\sum_{i=1}^{C}\hat{q}^{T}_{i}(x)\log q^{T}_{i}(x)$ (16)
-$=H\!\big(\hat{q}^{T}(x)\big)+\mathrm{KL}\!\big(\hat{q}^{T}(x)\|q^{T}(x)\big).$ (17)
-Where $H(\hat{q}^{T}(x))$ denotes the Shannon entropy of the teacher distribution,
-$H(\hat{q}^{T}(x)):=-\sum_{i=1}^{C}\hat{q}^{T}_{i}(x)\log\hat{q}^{T}_{i}(x),$
-so the KD loss decomposes into the teacher’s entropy and the KL divergence between the teacher and the student.
-The empirical distillation loss over $D_{\exp}$ is
-$\bar{L}_{\mathrm{KD}}=\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\mathcal{L}_{\mathrm{KD}}(x).$ (18)
-It serves as a measurable proxy for how closely the student imitates the teacher on the subset of old-task samples preserved through herding replay.
-Our goal is to study how much the intermediate-layer features drift after learning on a new task. For any layer $\ell$, define the drift as:
-$\Delta h_{\ell}(x)=h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1}).$ (19)
-which quantifies the amount by which the feature extractor changes its representation of an old-task sample.
-Bounding this is crucial for ensuring that LoRA plugins trained on earlier feature distributions remain effective after the model is updated on new tasks.
-In addition, to compare logits in a way that reflects changes in their meaningful structure rather than arbitrary additive shifts, we use centered logits:
-$\bar{z}(x)=z(x)-\tfrac{1}{C}\sum_{k=1}^{C}z_{k}(x)\mathbf{1},$ (20)
-$\bar{\hat{z}}(x)=\hat{z}(x)-\tfrac{1}{C}\sum_{k=1}^{C}\hat{z}_{k}(x)\mathbf{1}.$ (21)
-Centered logit is necessary because logits are defined only up to a constant offset: adding a constant to all logit entries leaves the softmax distribution unchanged.
-By removing this offset, we ensure that the norms of logit differences reflect genuine changes in class-relative predictions rather than uninformative bias shifts.
-This step allows the later Lipschitz bounds to connect logit discrepancies with feature discrepancies in a well-defined way.
-Assumptions
-A1 (Non-degenerate teacher distribution). There exists $\delta\in(0,1/C]$ that satisfies:
-$\hat{q}^{T}_{i}(x)\geq\delta\quad\text{for all }x\in D_{\exp}\text{ and }i.$ (22)
-This condition prevents extreme logit sharpness and ensures the log-softmax function is well-conditioned.
-The assumption is directly satisfied in standard knowledge distillation settings.
-In particular, the use of a temperature parameter $T>1$ produces softened teacher probabilities,
-$\hat{q}^{T}_{i}(x)=\frac{\exp(\hat{z}_{i}(x)/T)}{\sum_{j}\exp(\hat{z}_{j}(x)/T)},$
-which guarantees that no class probability collapses to zero.
-Alternatively, label smoothing also enforces a uniform lower bound on class probabilities.
-Thus, A1 follows from standard KD practice and is not a restrictive condition.
-A2 (Local bi-Lipschitz continuity).
-Let $\Phi_{\ell}$ denote the mapping from the $\ell$-th layer feature to the centered logits $\bar{z}$ through subsequent layers.
-There exist constants $0<\mu_{\ell}\leq L_{\ell}<\infty$ such that
-$\mu_{\ell}\|u-v\|_{2}\leq\|\Phi_{\ell}(u)-\Phi_{\ell}(v)\|_{2}\leq L_{\ell}\|u-v\|_{2},$ (23)
-for all $u,v$ within a neighborhood of past-task features in $D_{\exp}$.
-This ensures logits and features change proportionally, which is well supported by the structure of modern neural networks.
-Convolutional layers, affine transformations, and common nonlinearities (ReLU, SiLU, GELU) are all Lipschitz continuous.
-Since compositions of Lipschitz functions remain Lipschitz, the entire feature-to-logit map possesses an upper Lipschitz constant.
-The lower Lipschitz constant reflects local non-collapse of the representation: empirically, trained deep networks do not map distinct inputs from the same manifold region to identical logits.
-Similar assumptions have been used in analyses of embedding stability and invertibility of deep models.
-Therefore, A2 expresses a standard geometric regularity of neural networks rather than a special structural requirement.
-A3 (Empirical KD value).
-At the end of training, the empirical KD loss over $D_{\exp}$ equals $\bar{L}_{\mathrm{KD}}$.
-This is not an additional assumption but rather a definition of the observable optimization outcome.
-The analysis treats $\bar{L}_{\mathrm{KD}}$ as a known scalar summarizing how well the student matches the teacher on replay samples.
-A4 (Generalization from herded exemplars).
-The replay buffer $D_{\exp}$ maintains an equal number of samples per class, and $\mathcal{L}_{\mathrm{KD}}(x)\in[0,\log C]$ is bounded, so with probability at least $1-\delta_{0}$, we have:
-$\big|\mathbb{E}_{x\sim\mathcal{D}_{<t}}[\mathcal{L}_{\mathrm{KD}}(x)]-\bar{L}_{\mathrm{KD}}\big|\leq\epsilon_{\mathrm{gen}},$ (24)
-$\epsilon_{\mathrm{gen}}=O\!\Big(\sqrt{\tfrac{\log(1/\delta_{0})}{|D_{\exp}|}}\Big),$ (25)
-where $\mathcal{D}_{<t}$ is the joint distribution of all previous tasks.
-This assumption holds because the following properties guarantee that the empirical distillation loss concentrates around its population expectation.
-The boundedness of the distillation loss,
-$0\leq\mathcal{L}_{\mathrm{KD}}(x)\leq\log C,$
-implies that empirical averages of $\mathcal{L}_{\mathrm{KD}}$ satisfy standard concentration inequalities.
-Moreover, the herding procedure constructs $D_{\exp}$ as a class-balanced set that approximates class means in feature space, making it an approximately representative subset of the past-task distribution $\mathcal{D}_{<t}$.
-As a result, the empirical quantity
-$\bar{L}_{\mathrm{KD}}=\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\mathcal{L}_{\mathrm{KD}}(x)$
-forms a low-variance estimator of the population expectation. Applying Hoeffding-type bounds for bounded losses yields, with probability at least $1-\delta_{0}$,
-$\big|\mathbb{E}_{x\sim\mathcal{D}_{<t}}[\mathcal{L}_{\mathrm{KD}}(x)]-\bar{L}_{\mathrm{KD}}\big|\;\leq\;O\!\left(\sqrt{\tfrac{\log(1/\delta_{0})}{|D_{\exp}|}}\right),$
-Auxiliary Lemmas
-Lemma 1 (KD controls probability discrepancy). For any $x$, by Pinsker’s inequality:
-$\|q^{T}(x)-\hat{q}^{T}(x)\|_{1}\leq\sqrt{2\,\mathrm{KL}(\hat{q}^{T}(x)\|q^{T}(x))}\leq\sqrt{2(\mathcal{L}_{\mathrm{KD}}(x)-H(\hat{q}^{T}(x)))}.$ (26)
-Averaging this equation over $D_{\exp}$ gives
-$\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\|q^{T}(x)-\hat{q}^{T}(x)\|_{1}\leq\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\tfrac{1}{|D_{\exp}|}\sum_{x}H(\hat{q}^{T}(x))\Big)}.$
-Lemma 2 (Probability discrepancy controls logit discrepancy). Softmax relates logits and probabilities by:
-$(z_{i}-z_{j})=T\log(\frac{q^{T}_{i}}{q^{T}_{j}}).$ (27)
-So under A1, for any $i\neq j$, Because of Lipschitz continuity of $\log p$ over a bounded domain, differences in log-probabilities are linearly bounded by differences in probabilities:
-$|\log q^{T}_{i}-\log\hat{q}^{T}_{i}|\leq\frac{T}{\delta}(|q^{T}_{i}-\hat{q}^{T}_{i}|).$ (28)
-So it has:
-$|(z_{i}-z_{j})-(\hat{z}_{i}-\hat{z}_{j})|=T\Big|\log\frac{q^{T}_{i}}{q^{T}_{j}}-\log\frac{\hat{q}^{T}_{i}}{\hat{q}^{T}_{j}}\Big|$ (29)
-$\leq\frac{T}{\delta}(|q^{T}_{i}-\hat{q}^{T}_{i}|+|q^{T}_{j}-\hat{q}^{T}_{j}|).$ (30)
-Therefore,
-$\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\leq\frac{T}{\delta}\|q^{T}(x)-\hat{q}^{T}(x)\|_{1}.$ (31)
-Averaging Eq.(31) over $D_{\exp}$ and combining it with Eq. (B) in Lemma 1 yields:
-$\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\leq\frac{T}{\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\tfrac{1}{|D_{\exp}|}\sum_{x}H(\hat{q}^{T}(x))\Big)}.$ (32)
-Lemma 3. Under A2, for any $x$, it has:
-$\mu_{\ell}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}\leq\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}$ (33)
-$\leq L_{\ell}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}.$ (34)
-Main Conclusions
-Theorem 1 (Empirical bound on feature drift). For any layer $\ell$,
-$\frac{1}{|D_{\exp}|}\sum_{x\in D_{\exp}}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}\leq\frac{T}{\mu_{\ell}\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\tfrac{1}{|D_{\exp}|}\sum_{x}H(\hat{q}^{T}(x))\Big)}.$ (35)
-Proof. This conclusion can be straightforwardly proved by combining the Eq.(33) in Lemma 3 (first averaging it over $D_{\exp}$ ) with the Eq.(32) of Lemma 2.
-Theorem 2 (Distributional bound over past-task distribution). Under A1–A4, with probability at least $1-\delta_{0}$,
-$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\|_{2}\leq\frac{T}{\mu_{\ell}\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\mathbb{E}_{x\sim\mathcal{D}_{<t}}[H(\hat{q}^{T}(x))]+\epsilon_{\mathrm{gen}}\Big)}.$ (36)
-Proof. The expectation form of Lemma 2 states that
-$\ \mathbb{E}_{x\sim\mathcal{D}_{<t}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\;\leq\;\frac{T}{\delta}\sqrt{2\Big(\mathbb{E}_{x\sim\mathcal{D}_{<t}}\mathcal{L}_{\mathrm{KD}}(x)-\mathbb{E}_{x\sim\mathcal{D}_{<t}}H(\hat{q}^{T}(x))\Big)}.$ (37)
-A4 ensures that the empirical distillation loss generalizes to the past-task distribution:
-$\Big|\mathbb{E}_{x\sim\mathcal{D}_{<t}}\mathcal{L}_{\mathrm{KD}}(x)-\bar{L}_{\mathrm{KD}}\Big|\;\leq\;\epsilon_{\mathrm{gen}}.$ (38)
-Hence,
-$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\mathcal{L}_{\mathrm{KD}}(x)\;\leq\;\bar{L}_{\mathrm{KD}}+\epsilon_{\mathrm{gen}}.$ (39)
-Substituting this into Eq.(37) yields:
-$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}\;\leq\;\frac{T}{\delta}\sqrt{2\Big(\bar{L}_{\mathrm{KD}}-\mathbb{E}_{x\sim\mathcal{D}_{<t}}H(\hat{q}^{T}(x))+\epsilon_{\mathrm{gen}}\Big)}.$ (40)
-Lemma 3 states that for every $x$,
-$\mu_{\ell}\big\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\big\|_{2}\;\leq\;\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}.$ (41)
-Taking expectation on both sides gives:
-$\mathbb{E}_{x\sim\mathcal{D}_{<t}}\big\|h_{\ell}(x;\theta_{t})-h_{\ell}(x;\theta_{t-1})\big\|_{2}\;\leq\;\frac{1}{\mu_{\ell}}\mathbb{E}_{x\sim\mathcal{D}_{<t}}\|\bar{z}(x)-\bar{\hat{z}}(x)\|_{2}.$ (42)
-Then the conclusion to be proved can be obtained by combining Eq.(40) with Eq.(42).
-Theorems 1–2 imply
-$\mathbb{E}\|\Delta h_{\ell}(x)\|\leq\underbrace{\tfrac{T}{\mu_{\ell}\delta}}_{\text{constant }K_{\ell}}\times\underbrace{\sqrt{2(\Delta_{\mathrm{KD}}+\epsilon_{\mathrm{gen}})}}_{\text{variable }\Gamma_{t}},$ (43)
-where $\Delta_{\mathrm{KD}}=\mathbb{E}[\mathcal{L}_{\mathrm{KD}}(x)-H(\hat{q}^{T}(x))]$ reflects the effectiveness of distillation on the data of previous tasks.
-A smaller $\Delta_{\mathrm{KD}}$ represents the predictive behavior of the student model that aligns with that of the teacher model on these tasks.
-$\epsilon_{\mathrm{gen}}=O(\sqrt{\log(1/\delta_{0})/|D_{\exp}|})$ decreases as more exemplars are stored.
-The constant $\tfrac{T}{\mu_{\ell}\delta}$ depends on network smoothness and temperature parameters.
-The $K_{\ell}$ captures the conditioning of the feature-to-logit map and the softness of the teacher distribution, while the $\Gamma_{t}$ measures how much logit drift the distillation and replay objective permits.
-This bound shows that when distillation is optimized and the exemplars approximate the old-task distribution well, the deep feature representations of past data remain stable even as the model learns new tasks.
-Therefore, even though $\phi(\boldsymbol{x})$ continuously updates its parameters, the output of each layer for previous tasks remains relatively stable.
-This allows the plugins trained on earlier tasks to stay effective, thereby enabling the deployment of DLC in replay- and distillation-based CIL methods.